### v1

# Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in ultra low-data regimes

Abstract
Machine Learning (ML) in low-data settings remains an underappreciated yet crucial problem. This challenge is pronounced in low-to-middle income countries where access to large datasets is often limited or even absent. Hence, data augmentation methods to increase the sample size of datasets needed for ML are key to unlocking the transformative potential of ML in data-deprived regions and domains. Unfortunately, the limited training set constrains traditional tabular synthetic data generators in their ability to generate a large and diverse augmented dataset needed for ML tasks. To address this technical challenge, we introduce CLLM, which leverages the prior knowledge of Large Language Models (LLMs) for data augmentation in the low-data regime. While diverse, not all the data generated by LLMs will help increase utility for a downstream task, as for any generative model. Consequently, we introduce a principled curation process, leveraging learning dynamics, coupled with confidence and uncertainty metrics, to obtain a high-quality dataset. Empirically, on multiple real-world datasets, we demonstrate the superior performance of LLMs in the low-data regime compared to conventional generators. We further show our curation mechanism improves the downstream performance for all generators, including LLMs. Additionally, we provide insights and understanding into the LLM generation and curation mechanism, shedding light on the features that enable them to output high-quality augmented datasets. CLLM  paves the way for wider usage of ML in data scarce domains and regions, by allying the strengths of LLMs with a robust data-centric approach.
1 Introduction
No data, No Machine Learning. Machine learning (ML) has transformed numerous industries, but its wider adoption is hindered by a pervasive roadblock: insufficient data. Specifically, the use of ML algorithms presumes the availability and access to large datasets for training, be it in the form of labeled or unlabeled data. Unfortunately, real-world domains are often data scarce: (i) in healthcare and finance, collecting annotations can be expensive or practically impossible; (ii) in developing and low-to-middle income countries (LMICs), digital infrastructure (such as electronic healthcare records (EHRs)) can be limited or nonexistent (Ade-Ibijola and Okonkwo, 2023; Asiedu et al., 2023; Owoyemi et al., 2020; Mollura et al., 2020; Alami et al., 2020; Ciecierski-Holmes et al., 2022) and (iii) within large datasets, there can be (ethnic) minorities that are underrepresented. This lack of data has serious consequences: to sideline these settings to the peripheries of ML advancements and prevent the development of accurate models. How can we build a reliable ML model in this low-data regime, where we have so few samples? Solving this problem is a major opportunity that would unlock the potential of ML across society, domains, and regions.
Aim. To address this important yet undervalued low-data problem, we aim to augment the small labeled dataset ($n<100$) with synthetic samples. We focus on tabular data, as defining augmentations is non-trivial and can easily result in nonsensical or invalid samples. Moreover, tabular domains like healthcare (of value in LMICs) are often where data scarcity is acute.
Figure 1: CLLM  uses a small dataset $D_{\mathrm{train}}$ and a frozen black-box LLM to generate a larger synthetic set $D_{\mathrm{syn}}$. The curator computes the learning dynamics of samples in $D_{\mathrm{syn}}$, assessing samples based on their aleatoric uncertainty and predictive confidence, then curates $D_{\mathrm{syn}}$ with the goal that a downstream model trained on the curated $D_{\mathrm{curated}}$ will have improved performance.
Related work. Data augmentation is a widely used and different approach to address data scarcity in tabular data contexts. Methods are either based on generative models (Ghosheh et al., 2023; Biswas et al., 2023; Wang and Pai, 2023; Machado et al., 2022; Tanaka and Aranha, 2019) such as GANs (Xu et al., 2019), VAEs (Xu et al., 2019), Normalizing Flows (Papamakarios et al., 2021), Score-based models (Kotelnikov et al., 2022; Kim et al., 2022), or alternatively traditional methods such as SMOTE (Chawla et al., 2002; Wang and Pai, 2023; Machado et al., 2022). However, in ultra low-data regimes ($n<100$), the training data may not describe the full data distribution well, despite it being i.i.d. draws. Consequently, this harms conventional methods since the augmented data may not be sufficiently diverse and accurate, restricting the generalizability of predictive models trained on such data. Tangentially, prior works have tackled data scarcity in the tabular setting via the lens of transfer learning, where prior knowledge can be transferred from a pretrained model (Levin et al., 2022; Jin and Ucar, 2023) or a knowledge graph (Margeloiu et al., 2022; Ruiz et al., 2023), which might not be available in all settings. Recent work has shown the potential of fine-tuning Large Language Models (LLMs) for tabular data generation (Borisov et al., 2023). While LLMs offer some degree of prior knowledge, there are two challenges in our setting. First, it is computationally expensive to fine-tune LLMs, while needing specialized hardware —luxuries often not available in LMICs, thereby limiting applicability in such settings. Second, fine-tuning often assumes a large number of samples. In our low-data setting it could lead to overfitting and low-quality generated samples, and hence poor downstream models—as we show for Borisov et al. (2023) in Sec. 3.
Curated LLMs. To address these challenges, we propose Curated LLM (CLLM). First, CLLM  leverages the in-context capabilities of LLMs for generation, thereby reducing the computational burden. We also posit for the low-data regime; the diverse pretraining corpus of LLMs carries valuable prior knowledge, which may offer more diversity in their generation compared to other conventional tabular generators. Of course, LLMs are not perfect. Balancing the utility of LLMs against the risk of noisy, irrelevant data is important for downstream performance, hence requiring systemic assessment of the generated data. In fact, this issue is vital for any generative model.
This motivates the second key aspect of CLLM, i.e. post-generation data curation. This addresses the overlooked aspect that not all of the synthetic samples are useful to downstream model performance, with some samples even harmful. We anchor our approach with ideas from learning theory that show the behavior of individual data samples during training, called learning dynamics, provides a salient signal about the value of samples to a learner (Arpit et al., 2017; Arora et al., 2019; Li et al., 2020). To provide intuition, samples with variable predictions might be considered ambiguous or other samples might never be learned correctly and could harm a model. In CLLM, we study the learning dynamics of the synthetic data samples, with respect to a model trained on the small real dataset. We then analyze these dynamics by computing two key metrics: confidence and aleatoric (data) uncertainty. These metrics form the basis for curating the synthetic samples. We aim to enable a highly performant downstream model when trained on the curated dataset.
Contributions: CLLM  is a novel data augmentation approach allying the strengths of LLMs with a robust data curation mechanism to improve data augmentation in the ultra low-data regime $(n<100)$, bringing several contributions: 1⃝ Improved performance: we empirically demonstrate on 7 real-world datasets that CLLM  enables superior downstream performance compared to 6 widely used tabular data generative models and data augmentation techniques. 2⃝ Value of curation: we show the overlooked aspect of synthetic data curation improves downstream performance across the generative models. This highlights the flexibility and broad utility of our curation mechanism for data augmentation. 3⃝ Insights: we dissect the two aspects of CLLM (LLM and data curation) along a variety of dimensions, providing insights and understanding into why the approach is beneficial. We show the largest gains are for underrepresented subgroups and in ultra low-data settings. These contributions pave the way towards wider usage of ML across society, domains and regions.
Ethical considerations. LLMs may make errors and may reflect or exacerbate societal biases that are present in their data (Li et al., 2023). Though the curation in CLLM improves synthetic data quality, it does not directly aim to remove biases. The quality and fairness of generated data should always be evaluated. More research into LLM bias is required before methods like CLLM should be applied to real-world sensitive settings like healthcare and finance.
2 CLLM: Synergy of LLM Generation and Data Curation
Set-up. Given feature space $\mathcal{X}$, and label space $\mathcal{Y}=\{1,...,k\}$, we assume that we only have a small labeled dataset $D_{\mathrm{train}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, with $x_{i}\in\mathcal{X}$, $y_{i}\in\mathcal{Y}$ and $n<100$ (ultra-low data setting). Assume $D_{\mathrm{train}}$ is drawn i.i.d. from the real distribution $p_{R}(X,Y)$. We also assume access to a pretrained LLM to generate samples. We denote the output distribution of the LLM as $p_{\Phi}(X,Y)$, with $\Phi$ containing parameters that we control (e.g., input prompts). Our goal is to generate a dataset to augment the small $D_{\mathrm{train}}$, and subsequently use it to train a classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$. Successful augmentation will provide a better classifier $f$, than if we had trained $f$ on the small $D_{\mathrm{train}}$ itself. We measure downstream performance on a separate held-out dataset of real data, $D_{\mathrm{test}}$.
Our Approach. To address this challenge, we introduce CLLM, an approach for data augmentation in low-data regimes. As shown in Figure 1, CLLM  leverages LLMs to generate a synthetic dataset $D_{\mathrm{syn}}$ using a small dataset $D_{\mathrm{train}}$ (Sec. 2.1). It exploits the LLMs’ prior knowledge via in-context learning (ICL) and contextual information. CLLM  then curates $D_{\mathrm{syn}}$ by analyzing the learning dynamics of samples in $D_{\mathrm{syn}}$ based on predictive confidence and aleatoric (data) uncertainty. These metrics are obtained by training a supervised model on $D_{\mathrm{train}}$. We leverage them to define a curated dataset $D_{\mathrm{curated}}$, which is used to train a downstream classifier (Sec. 2.2).
In each sub-section we describe and motivate the design of the different aspects of CLLM  (LLM and curation). Furthermore, we provide insights and understanding into their role in improving data utility, which we later quantify on multiple real-world datasets in Sec. 3.
2.1 Data generation with LLMs based on a small $D_{\mathrm{train}}$
As outlined in Sec. 1, in the ultra low-data regime, conventional tabular generative models (e.g. CTGAN, TVAE) are constrained by the limited $D_{\mathrm{train}}$ and may not generate sufficiently diverse and/or accurate synthetic data. To address this challenge, we propose to leverage LLMs, building on their large-scale pretraining. We first outline the desirable features of LLMs for tabular data generation when we have very few samples, then describe design choices to satisfy these.
  • Prior knowledge. LLMs have been pretrained with a vast corpus of information (Chowdhery et al., 2022; Singhal et al., 2023). When prompted to generate samples with limited real data, LLMs can leverage this encoded prior information about similar problems and feature-label relationships to enhance both accuracy and diversity of generation.
  • Contextual understanding. LLMs can process background and contextual information about the problem via natural language (Yang et al., 2023). For example, a high-level description of the task, features and their meanings can be conveniently described through natural language. Such information is unavailable to conventional generators that only utilize numerical examples.
  • Few-shot capabilities. LLMs have demonstrated proficiency in generalizing to tasks with just a few examples (Brown et al., 2020; Wei et al., 2023; Mirchandani et al., 2023). In the context of generation, we envision the idea of in-context generation using limited real examples.
To benefit from these capabilities, we craft the LLM prompt with three different parts (see Fig. 1): (1) Background: text description of the dataset and task (e.g. predict Covid mortality). Additionally, we include a description of what each feature means, explicitly prompting the LLM to use prior knowledge about these features. (2) Examples: we serialize the samples in $D_{\mathrm{train}}$ as example demonstrations and provide both the features and the label in text format. (3) Instructions: To generate a synthetic dataset $D_{\mathrm{syn}}$, we instruct the LLM to leverage the contextual information and provided examples as an i.i.d. draw from the distribution. We instruct the LLM to identify structural and feature-label relationships in the data and generate diverse data following the structure and format of the provided examples. We provide more details on the prompts in Appendix B.
Motivation for a frozen LLM. Using a frozen black-box LLM (e.g. GPT-4 or GPT-3.5) is computationally cheaper and requires less specialized hardware (i.e. GPUs) compared to fine-tuning. This relates to settings described in Sec. 1, such as LMICs, where we may not have the computational resources to fine-tune an LLM. Even in settings where fine-tuning is possible, we show empirically in Sec. 3 that LLM fine-tuning (e.g. GReaT baseline) is suboptimal in ultra-low data settings ($n<100$) compared to providing in-context examples coupled with curation.
Dissecting the LLM’s generative features. We now investigate various dimensions to understand and illustrate empirically the appealing features of LLMs as data generators in the low-data regime, and how our design choices unlock them. We take the Brazilian Covid-19 dataset (Baqui et al., 2020) as a running example and focus on GPT-4 as the LLM.
[…]
Figure 2: GPT-4 is able to extrapolate to regions of the oracle (true manifold) even where there is no training data covering them, as can be seen by the overlap with the turquoise dots, with the effect more pronounced when $D_{\mathrm{train}}$ is small
[…]
[Uncaptioned image] Figure 3: Subgroups with fewest samples in $D_{\mathrm{train}}$ benefit the most from data augmentation, on average. Table 1: Deep dive into the top 5 demographic subgroups in the Covid dataset with the largest gains, across 10 seeds, for $|D_{\mathrm{train}}|=20$. GPT-4 improves performance on the smallest groups. Subgroup $n_{\mathrm{samples}}$ in $D_{\mathrm{train}}$ Avg. Acc. Gain v. $D_{\mathrm{train}}$ (min - max) GPT-4 TVAE Age_40 0-6 6.38 +- 2.09 -3.37 +- 2.86 Liver 0-1 3.85 +- 3.37 -13.1 +- 3.38 Renal 0-3 4.52 +- 2.01 -18.0 +- 3.22 Amarela 0-1 8.71 +- 1.40 -2.03 +- 2.88 Parda 3-11 5.07 +- 1.50 -6.57 +- 1.61
[…]
[Uncaptioned image] Figure 4: Contextual information in the prompt is important for extrapolation. Table 2: Including contextual information in the prompt improves precision (P), recall (R), and utility (U) in low-sample settings (results shown for the Covid dataset). $n_{\mathrm{samples}}$ in $D_{\mathrm{train}}$ GPT-4 w/ context GPT-4 no context TVAE P R U P R U P R U 20 $\bf 0.41_{(0.04)}$ $\bf 0.87_{(0.03)}$ $\bf 0.74_{(0.01)}$ $0.13_{(0.0)}$ $0.82_{(0.01)}$ $0.66_{(0.01)}$ $0.33_{(0.07)}$ $0.50_{(0.03)}$ $0.59_{(0.02)}$ 40 $\bf 0.40_{(0.01)}$ $\bf 0.91_{(0.01)}$ $\bf 0.76_{(0.0)}$ $0.11_{(0.0)}$ $0.89_{(0.0)}$ $0.69_{(0.0)}$ $0.27_{(0.01)}$ $0.68_{(0.01)}$ $0.62_{(0.03)}$ 100 $\bf 0.42_{(0.01)}$ $0.86_{(0.02)}$ $\bf 0.75_{(0.01)}$ $0.11_{(0.01)}$ $\bf 0.90_{(0.01)}$ $0.74_{(0.01)}$ $0.39_{(0.02)}$ $0.67_{(0.03)}$ $0.64_{(0.06)}$ 200 $0.44_{(0.02)}$ $0.85_{(0.02)}$ $\bf 0.75_{(0.0)}$ $0.08_{(0.01)}$ $\bf 0.90_{(0.0)}$ $0.60_{(0.01)}$ $\bf 0.47_{(0.0)}$ $0.73_{(0.01)}$ $0.65_{(0.02)}$
2.2 Data curation with learning dynamics
When prompted with $\Phi$ (which contains the in-context samples of $D_{\mathrm{train}}$), the LLM generates samples from a distribution $p_{\Phi}(X,Y)$ that approximates $p_{R}(X,Y)$, implicitly exploiting its large-scale pretraining and few-shot capabilities. LLMs are of course not perfect and could generate noisy samples, hence this distribution may be inaccurate  [^1] We could finetune the model on the scarce $D_{\mathrm{train}}$ we have, but is likely to still lead to overfitting due to the extreme data scarcity and LLM parameter size.. To make this distribution more relevant to the downstream task, we include a data curation mechanism. Specifically, we focus on the noisy feature-label relationship $p_{\Phi}(Y|X)$, for which we expect $p_{\Phi}(Y|X)\neq p_{R}(Y|X)$ given the small size of $D_{\mathrm{train}}$. This motivates us to curate $D_{\mathrm{syn}}$ and discard likely mislabeled samples.
We anchor our approach with ideas from learning theory that show the behavior of individual samples during model training (called learning dynamics) contains signal about the nature of the samples themselves (Arpit et al., 2017; Arora et al., 2019; Li et al., 2020). Some samples are easily and confidently predicted over different model checkpoints, whereas other samples might be challenging (e.g. due to mislabeling) and hence might be incorrectly predicted for the given label. Consequently, we operationalize learning dynamics as the basis of our curation mechanism. Specifically, we analyze samples in $D_{\mathrm{syn}}$ by studying their learning dynamics computed with a classifier trained on $D_{\mathrm{train}}$. We then categorize and filter samples in $D_{\mathrm{syn}}$, and produce a curated dataset $D_{\mathrm{curated}}\subset D_{\mathrm{syn}}$.
Learning dynamics. We now formalize how we compute learning dynamics for individual samples. Assume that a classifier $f$ is trained in an iterative scheme (e.g. neural networks or XGBoost trained over iterations) on $D_{\mathrm{train}}$, which makes it possible to analyze the learning dynamics of samples in $D_{\mathrm{syn}}$ over these iterations. The classifier $f$ should be at least as flexible as the model that the practitioner intends to use for the downstream task. $f$ is trained from scratch on $D_{\mathrm{train}}$ and goes through $e\in[E]$ different checkpoints leading to the set =$\mathcal{F}=\{f_{1},f_{2},\dots,f_{E}\}$, such that $f_{e}$ is the classifier at the $e$-th checkpoint. Let $[f_{e}(x)]_{y}$ denote the predicted probability for class $y$ and sample $x$. Our goal is to assess the learning dynamics of samples in $D_{\mathrm{syn}}$ over these $E$ training checkpoints, while we train $f$ on $D_{\mathrm{train}}$. For this, we define $H$, a random variable following a uniform distribution $\mathcal{U}_{\mathcal{F}}$ over the set of checkpoints $\mathcal{F}$. Specifically, given $H=h$ and a sample $(x,y)$, we define the correctness in the prediction of $H$ as a binary random variable $\hat{Y}_{\mathcal{F}}(x,y)$ with the following conditional: $P(\hat{Y}_{\mathcal{F}}(x,y)=1|H=h)=[h(x)]_{y}$ and $P(\hat{Y}_{\mathcal{F}}(x,y)=0|H=h)=1-P(\hat{Y}_{\mathcal{F}}(x,y)=1|H=h)$.
Curation metrics. Equipped with a probabilistic interpretation of the predictions of a model, we now define two characterization metrics that we use for curation: (i) average confidence and (ii) aleatoric (data) uncertainty, inspired by (Kwon et al., 2020; Seedat et al., 2022a).
Definition 2.1 (Average confidence).
For any set of checkpoints =$\mathcal{F}=\{f_{1},...,f_{E}\}$, the average confidence for a sample $(x,y)$ is defined as the following marginal:
$\bar{P}_{\mathcal{F}}(x,y):=P(\hat{Y}_{\mathcal{F}}(x,y)=1)=\mathbb{E}_{H\sim\mathcal{U}_{\mathcal{F}}}[P(\hat{Y}_{\mathcal{F}}(x,y)=1|H)]=\frac{1}{E}\sum_{e=1}^{E}[f_{e}(x)]_{y}$ (1)
Definition 2.2 (Aleatoric uncertainty).
For any set of checkpoints =$\mathcal{F}=\{f_{1},...,f_{E}\}$, the aleatoric uncertainty for a sample $(x,y)$ is defined as:
$v_{al,\mathcal{F}}(x,y):=\mathbb{E}_{H\sim\mathcal{U}_{\mathcal{F}}}[Var(\hat{Y}_{\mathcal{F}}(x,y)|H)]=\frac{1}{E}\sum_{e=1}^{E}[f_{e}(x)]_{y}(1-[f_{e}(x)]_{y})$ (2)
Intuitively, for binary classification ($k=2$), the aleatoric uncertainty for a sample $x$ is maximized when $[f_{e}(x)]_{y}=\frac{1}{2}$ for all checkpoints $f_{e}$, akin to random guessing. Recall aleatoric uncertainty captures the inherent data uncertainty, hence is a principled way to capture issues such as mislabeling. This contrasts epistemic uncertainty, which is model-dependent and can be reduced simply by increasing model parameterization (Hüllermeier and Waegeman, 2021).
Having defined sample-wise confidence and aleatoric uncertainty, we characterize samples in $D_{\mathrm{syn}}$ into two categories, namely Selected and Discarded. Given a sample $(x,y)$, a set of training checkpoints $\mathcal{F}$, and two thresholds $\tau_{\mathrm{conf}}$ and $\tau_{\mathrm{al}}$, we define the category $c(x,y,\mathcal{F})$ as Discarded if $\bar{P}_{\mathcal{F}}(x,y)<\tau_{\mathrm{conf}}$ and $v_{al,\mathcal{F}}(x,y)<\tau_{\mathrm{al}}$, and Selected otherwise.
[…]
To summarize, the objective of the curation step is that training on the curated synthetic data leads to a better classifier $f_{D_{\mathrm{curated}}}$ for the downstream task, compared to training on the uncurated synthetic data, i.e. $M(f_{D_{\mathrm{curated}}})>M(f_{\mathcal{D_{\mathrm{syn}}}})$, where $M$ is a performance measure (for example accuracy). In Sec. 3, we empirically show how performance on this curated dataset is superior both for LLM generated data as well as other classes of generative models.
Dissecting the role of curation. We now empirically demonstrate the role of curation in correcting the noisy feature-label relationship present in $D_{\mathrm{syn}}$, highlighting two insights: (i) curation discards samples which are atypical in their label with respect to their neighbors in $D_{\mathrm{syn}}$ (ii) discarded samples can be considered “mislabeled”, and we quantify their atypicality using a large held-out dataset $D_{\mathrm{oracle}}$.
[…]
We highlight two key observations. First, the curated datasets, for all the generative models, exhibit a higher agreement with the proxy $\hat{\eta}$ than the discarded datasets. This aligns with the desideratum of only keeping samples that exhibit the correct feature-label relationships.
Figure 5: $\hat{\eta}$ aligns more with $D_{\mathrm{curated}}$ than $D_{\mathrm{discarded}}$ for each generative model: the curation step keeps high quality samples tailored to the downstream task.
This provides a rationale for why curation helps improve discriminative performance, as samples in $D_{\mathrm{curated}}$ are much more likely to have the correct feature-label relationship.
Second, GPT-4 has a higher agreement with $\hat{\eta}$ on $D_{\mathrm{discarded}}$, compared to the other generators. This illustrates that GPT-4’s prior knowledge enables it to better capture the distribution $p(Y|X=x)$. Note that generative baselines (e.g. TVAE) model the joint $p(X,Y)$, without any context of which is the set of features and which is the label. In contrast, we can define in the LLM prompt which column is the target $Y$, allowing the LLM to better capture the feature-label relationships. This complements the findings from Fig. 2, which showed that GPT-4 extrapolates to unseen regions of the feature manifold, captured by the support of $p(X)$.
3 Curated LLMs for Better Data Augmentation
We now perform an end-to-end quantitative evaluation of CLLM  across multiple real-world datasets, for downstream utility, demonstrating the value of allying the generative capabilities of LLMs with our curation mechanism. Sec. 3.1 compares the performance of GPT-4 and our curation approach with respect to a variety of state-of-the-art tabular augmentation baselines. Having evaluated CLLM  on a range of datasets, we also demonstrate how we can leverage information extracted during curation to characterize datasets via a hardness proxy. Sec. 3.2 illustrates how our characterization of samples during the curation step can help to flag synthesized datasets (e.g via the LLM) which, if used for training, will result in poor downstream performance.
Experimental setup. We compare CLLM  (with GPT-4 (OpenAI, 2023) and GPT-3.5 (Brown et al., 2020)) against a variety of baselines for tabular data generation and augmentation: CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), Normalizing Flows (Papamakarios et al., 2021), TabDDPM (Kotelnikov et al., 2022), SMOTE (Chawla et al., 2002) and GReaT (Borisov et al., 2023), which fine-tunes an LLM. We evaluate performance on 7 real-world datasets with different feature counts and vary the number of samples available in $D_{\mathrm{train}}$, repeating each experiment across $10$ seeds.
While we do not know the exact makeup of the pretraining data of LLMs like GPT-4, there is the possibility that open-source data might be included. This poses a risk of memorization as the primary source of performance gain. To disentangle the role of memorization, we select 4 real-world medical datasets (Maggic (Pocock et al., 2013), Covid (Baqui et al., 2020), SEER (Duggan et al., 2016), CUTRACT (PCUK, 2019)) that require an authorization process to access, hence are unlikely to form part of the LLMs training corpus. We use common open-source datasets (Adult and Drug from the UCI repository (Asuncion and Newman, 2007) and Compas (Angwin et al., 2016)) that are highly reflective of data scarce domains. Further experimental details can be found in Appendix B.
Table 3: AUC averaged over 4 downstream models on $D_{\mathrm{test}}$ where curation improves performance for all methods across all sample sizes $n$, as indicated by $\uparrow$. CLLM  w/ GPT-4 (Curated) dataset provides the strongest performance for both private/proprietary datasets and public datasets
Real data CLLM (OURS) Baselines
GPT-4 GPT-3.5 CTGAN TabDDPM GReaT NFLOW SMOTE TVAE
Dataset $D_{\mathrm{oracle}}$ $D_{\mathrm{train}}$ Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur.
covid (n=20) 74.41 68.50 73.78 73.87 normal-$\uparrow$ 69.85 71.41 $\uparrow$ 59.00 63.67 $\uparrow$ 66.84 66.85 $\uparrow$ 57.38 66.46 $\uparrow$ 62.87 68.56 $\uparrow$ 66.95 66.82 61.69 66.11 $\uparrow$
cutract (n=20) 72.23 70.12 71.15 72.50 normal-$\uparrow$ 69.97 71.54 $\uparrow$ 64.01 67.98 $\uparrow$ 66.05 66.59 $\uparrow$ 52.38 67.02 $\uparrow$ 64.44 70.42 $\uparrow$ 68.41 69.24 $\uparrow$ 68.94 70.22 $\uparrow$
maggic (n=20) 67.41 57.13 60.70 61.48 normal-$\uparrow$ 57.54 58.69 $\uparrow$ 52.75 54.51 $\uparrow$ 54.59 55.39 $\uparrow$ 50.29 55.64 $\uparrow$ 54.72 57.38 $\uparrow$ 55.84 56.15 $\uparrow$ 54.08 56.19 $\uparrow$
seer (n=20) 87.92 80.67 84.53 84.82 normal-$\uparrow$ 83.34 83.71 $\uparrow$ 74.34 78.73 $\uparrow$ 80.59 80.60 $\uparrow$ 47.57 74.43 $\uparrow$ 76.06 79.98 $\uparrow$ 79.23 80.02 $\uparrow$ 74.53 78.73 $\uparrow$
compas (n=20) 67.51 63.11 68.01 67.91 62.07 64.43 $\uparrow$ 55.67 62.56 $\uparrow$ 57.67 60.87 $\uparrow$ 53.33 63.59 $\uparrow$ 59.49 64.62 $\uparrow$ 61.06 61.59 $\uparrow$ 58.30 62.58 $\uparrow$
adult (n=20) 84.17 77.45 50.39 71.48 $\uparrow$ 49.23 72.37 $\uparrow$ 72.23 76.86 $\uparrow$ 74.35 75.04 $\uparrow$ 67.00 77.25 $\uparrow$ 67.46 76.48 $\uparrow$ 73.75 73.67 73.20 76.90 $\uparrow$
drug (n=20) 77.81 70.84 75.08 75.29 normal-$\uparrow$ 71.68 72.14 $\uparrow$ 68.31 72.65 $\uparrow$ 68.12 69.68 $\uparrow$ 58.78 68.89 $\uparrow$ 62.13 67.75 $\uparrow$ 70.16 70.16 66.60 69.18 $\uparrow$
covid (n=40) 75.02 70.77 73.40 73.95 normal-$\uparrow$ 70.42 71.93 $\uparrow$ 63.63 68.46 $\uparrow$ 70.50 70.44 56.50 68.68 $\uparrow$ 66.41 70.48 $\uparrow$ 68.66 68.44 61.03 67.35 $\uparrow$
cutract (n=40) 72.57 69.18 69.87 71.72 normal-$\uparrow$ 68.47 69.56 $\uparrow$ 63.01 67.87 $\uparrow$ 65.63 67.27 $\uparrow$ 54.39 68.44 $\uparrow$ 61.40 67.98 $\uparrow$ 67.86 67.95 $\uparrow$ 59.79 66.62 $\uparrow$
maggic (n=40) 67.50 58.26 59.29 60.77 normal-$\uparrow$ 57.50 59.15 $\uparrow$ 55.00 56.78 $\uparrow$ 55.24 56.94 $\uparrow$ 48.81 56.64 $\uparrow$ 54.68 58.58 $\uparrow$ 57.40 57.44 $\uparrow$ 55.04 57.33 $\uparrow$
seer (n=40) 87.90 82.93 84.29 84.93 normal-$\uparrow$ 83.46 84.44 $\uparrow$ 80.05 83.67 $\uparrow$ 82.59 81.37 54.93 81.11 $\uparrow$ 79.88 84.36 $\uparrow$ 80.79 82.21 $\uparrow$ 78.69 83.62 $\uparrow$
compas (n=40) 67.35 62.34 67.57 67.85 normal-$\uparrow$ 61.34 62.84 $\uparrow$ 56.29 61.02 $\uparrow$ 58.85 60.11 $\uparrow$ 58.88 64.37 $\uparrow$ 58.61 63.54 $\uparrow$ 60.83 60.95 $\uparrow$ 55.94 61.04 $\uparrow$
adult (n=40) 84.43 79.44 48.31 73.82 $\uparrow$ 49.21 74.27 $\uparrow$ 71.82 79.11 $\uparrow$ 71.51 77.99 $\uparrow$ 66.77 78.81 $\uparrow$ 71.13 79.71 $\uparrow$ 77.90 78.84 $\uparrow$ 72.58 80.02 normal-$\uparrow$
drug (n=40) 77.71 71.86 74.30 75.79 normal-$\uparrow$ 71.33 72.76 $\uparrow$ 69.46 72.74 $\uparrow$ 71.08 73.07 $\uparrow$ 64.89 73.64 $\uparrow$ 62.51 70.97 $\uparrow$ 69.23 69.78 $\uparrow$ 65.22 70.30 $\uparrow$
covid (n=100) 74.52 71.57 73.77 74.71 normal-$\uparrow$ 70.71 72.76 $\uparrow$ 69.05 72.13 $\uparrow$ 71.60 73.22 $\uparrow$ 63.52 72.04 $\uparrow$ 64.25 72.64 $\uparrow$ 70.08 70.78 $\uparrow$ 69.05 71.96 $\uparrow$
cutract (n=100) 72.36 70.96 70.20 72.51 normal-$\uparrow$ 69.97 71.94 $\uparrow$ 67.94 72.42 $\uparrow$ 70.53 71.98 $\uparrow$ 55.72 69.14 $\uparrow$ 67.59 72.42 $\uparrow$ 68.79 69.68 $\uparrow$ 66.89 71.52 $\uparrow$
maggic (n=100) 67.46 59.65 58.98 61.32 normal-$\uparrow$ 55.71 58.90 $\uparrow$ 57.20 59.34 $\uparrow$ 57.26 58.28 $\uparrow$ 49.54 57.91 $\uparrow$ 56.36 60.11 $\uparrow$ 58.89 58.99 $\uparrow$ 56.17 58.86 $\uparrow$
seer (n=100) 87.79 83.95 84.45 85.37 normal-$\uparrow$ 83.92 85.08 $\uparrow$ 81.60 85.14 $\uparrow$ 83.04 84.83 $\uparrow$ 70.32 83.83 $\uparrow$ 81.16 85.03 $\uparrow$ 81.82 82.49 $\uparrow$ 78.88 84.50 $\uparrow$
compas (n=100) 67.18 62.56 68.02 68.19 normal-$\uparrow$ 60.10 62.47 $\uparrow$ 60.01 63.73 $\uparrow$ 58.32 61.34 $\uparrow$ 59.97 64.19 $\uparrow$ 60.02 64.04 $\uparrow$ 61.44 61.73 $\uparrow$ 59.97 62.82 $\uparrow$
adult (n=100) 84.34 81.24 46.09 74.57 $\uparrow$ 47.56 73.97 $\uparrow$ 74.29 80.45 $\uparrow$ 75.93 78.22 $\uparrow$ 77.09 81.66 normal-$\uparrow$ 70.70 81.04 $\uparrow$ 80.56 81.10 $\uparrow$ 74.04 80.23 $\uparrow$
drug (n=100) 78.00 73.58 76.24 76.74 normal-$\uparrow$ 69.46 71.05 $\uparrow$ 68.19 73.28 $\uparrow$ 72.43 73.79 $\uparrow$ 67.26 75.28 $\uparrow$ 62.67 73.12 $\uparrow$ 70.90 71.53 $\uparrow$ 68.22 73.59 $\uparrow$
covid (n=200) 74.69 72.33 73.40 74.62 normal-$\uparrow$ 70.70 73.12 $\uparrow$ 71.07 73.89 $\uparrow$ 72.47 74.44 $\uparrow$ 65.55 73.07 $\uparrow$ 65.04 72.90 $\uparrow$ 71.68 71.87 $\uparrow$ 67.89 72.38 $\uparrow$
cutract (n=200) 72.52 71.75 71.39 73.01 $\uparrow$ 70.28 72.39 $\uparrow$ 69.28 72.41 $\uparrow$ 71.83 74.03 normal-$\uparrow$ 66.66 72.49 $\uparrow$ 68.77 73.16 $\uparrow$ 70.23 70.80 $\uparrow$ 66.61 71.87 $\uparrow$
maggic (n=200) 67.37 61.39 58.92 61.41 normal-$\uparrow$ 57.33 60.16 $\uparrow$ 58.48 61.33 $\uparrow$ 56.26 57.20 $\uparrow$ 50.74 59.60 $\uparrow$ 55.95 60.75 $\uparrow$ 60.73 60.78 $\uparrow$ 57.18 60.23 $\uparrow$
seer (n=200) 87.84 84.63 84.39 85.56 $\uparrow$ 83.48 84.80 $\uparrow$ 82.04 85.34 $\uparrow$ 84.39 86.57 normal-$\uparrow$ 82.15 86.03 $\uparrow$ 77.73 85.19 $\uparrow$ 83.38 84.15 $\uparrow$ 79.71 85.26 $\uparrow$
compas (n=200) 67.14 63.27 67.02 68.15 normal-$\uparrow$ 60.48 63.39 $\uparrow$ 60.58 64.32 $\uparrow$ 60.60 63.52 $\uparrow$ 61.11 65.08 $\uparrow$ 56.58 63.60 $\uparrow$ 61.99 62.80 $\uparrow$ 60.15 63.99 $\uparrow$
adult (n=200) 84.25 82.12 40.96 75.84 $\uparrow$ 49.89 76.11 $\uparrow$ 78.18 82.32 $\uparrow$ 81.66 83.17 $\uparrow$ 80.06 83.32 normal-$\uparrow$ 74.31 82.64 $\uparrow$ 82.26 82.39 $\uparrow$ 75.21 82.02 $\uparrow$
drug (n=200) 77.36 76.10 75.58 76.06 $\uparrow$ 70.66 72.81 $\uparrow$ 71.31 75.98 $\uparrow$ 69.61 71.79 $\uparrow$ 72.35 77.41 normal-$\uparrow$ 65.25 75.26 $\uparrow$ 74.38 74.78 $\uparrow$ 68.39 74.33 $\uparrow$
3.1 Overall performance: downstream utility
We assess overall performance based on Utility of the augmented data, which we evaluate in terms of AUC on the real $D_{\mathrm{test}}$, when using four different types of downstream models (see Appendix B). This setup mirrors the widely adopted Train-on-synthetic-Test-on-real (TSTR) (Esteban et al., 2017). Additionally, we compare the performance to training on the small $D_{\mathrm{train}}$, as well as training on the large held-out $D_{\mathrm{oracle}}$, the latter serving as an upper bound.
GPT-4 + Curation has best overall performance. Table 3 shows the performance of the proposed CLLM  (GPT- 4 and GPT-3.5) against baselines — both with and without our curation mechanism. We find that the GPT-4 + Curation variant of CLLM  outperforms baselines in almost all settings ($20/28$). Interestingly, its performance is close to or even exceeds the performance of $D_{\mathrm{oracle}}$. Table 4 further shows that GPT-4 + Curation ranks first on average among all the generative methods.
Sample size sensitivity. We now investigate the performance gains of CLLM  as we vary the number of samples $n$ in $D_{\mathrm{train}}$, in Table 3 and Table 4. Performance improvements and high ranking across datasets for CLLM (GPT-4+Curation) are especially noticeable in the ultra low-data regime (i.e. $n<100$). In this regime, the limited size of $D_{\mathrm{train}}$ severely constrains the other baseline methods. In contrast, as illustrated in Sec. 2.1, CLLM  can leverage GPT-4’s prior knowledge to extrapolate beyond the small $D_{\mathrm{train}}$, thereby improving downstream performance. As expected, the performance gap between CLLM  and other methods decreases as the size of $D_{\mathrm{train}}$ grows (e.g. $n=200$), where sufficient training data helps other generators achieve good performance. We further decouple the prior knowledge of the LLM and the number of in-context samples in Appendix C.3, showing the importance of the in-context samples to guide the LLM’s generation.
Table 4: Average rank of approaches across the different datasets and seeds. CLLM  w/ GPT-4 ranks first across all $n$ and curation improves all the generative models.
Method n=20 n=40 n=100 n=200
CLLM  w/ GPT-4 2.71 $\pm$ 1.44 2.14 $\pm$ 1.06 2.29 $\pm$ 1.19 3.29 $\pm$ 1.38
GPT-4 3.86 $\pm$ 1.73 4.29 $\pm$ 1.83 6.00 $\pm$ 1.77 7.57 $\pm$ 1.65
CLLM  w/ GPT-3.5 4.14 $\pm$ 0.94 4.14 $\pm$ 0.71 6.86 $\pm$ 1.24 7.57 $\pm$ 0.70
NFLOW (curated) 6.00 $\pm$ 1.21 4.71 $\pm$ 0.80 4.00 $\pm$ 0.57 4.71 $\pm$ 0.63
GPT-3.5 6.71 $\pm$ 1.52 7.29 $\pm$ 1.26 11.57 $\pm$ 0.94 12.57 $\pm$ 0.57
TVAE (curated) 7.14 $\pm$ 1.17 7.86 $\pm$ 1.30 6.43 $\pm$ 0.40 6.71 $\pm$ 0.52
SMOTE (curated) 7.71 $\pm$ 0.33 8.14 $\pm$ 0.91 7.71 $\pm$ 1.19 7.43 $\pm$ 1.07
SMOTE 7.86 $\pm$ 0.55 9.57 $\pm$ 0.80 9.57 $\pm$ 1.09 9.00 $\pm$ 1.03
TabDDPM (curated) 8.29 $\pm$ 0.98 8.00 $\pm$ 0.93 6.00 $\pm$ 0.95 5.14 $\pm$ 1.68
CTGAN (curated) 8.29 $\pm$ 1.42 7.14 $\pm$ 0.91 4.14 $\pm$ 0.62 3.71 $\pm$ 0.39
GReaT (curated) 8.57 $\pm$ 1.50 6.57 $\pm$ 1.21 6.29 $\pm$ 1.38 3.57 $\pm$ 0.92
TabDDPM 10.14 $\pm$ 1.19 9.86 $\pm$ 1.15 10.00 $\pm$ 1.03 10.29 $\pm$ 1.02
TVAE 12.14 $\pm$ 0.89 14.00 $\pm$ 0.70 13.71 $\pm$ 0.39 14.43 $\pm$ 0.40
NFLOW 12.86 $\pm$ 0.47 14.14 $\pm$ 0.37 14.00 $\pm$ 0.45 15.29 $\pm$ 0.33
CTGAN 13.86 $\pm$ 0.68 13.14 $\pm$ 0.47 12.86 $\pm$ 0.37 12.00 $\pm$ 0.53
GReaT 15.71 $\pm$ 0.26 15.00 $\pm$ 0.53 14.57 $\pm$ 1.03 12.71 $\pm$ 0.96
Curation generally helps all generative models. Our curation mechanism consistently benefits all generative models for the different $n$. It ensures only high quality samples are retained, which is crucial for good data augmentation and downstream performance and has been overlooked in previous works. This explains why the combination of the best generative model and curation, which is CLLM, gives the best results and highest rankings in the low-data regime (e.g. $n=20$).
Performance benefits maintained for private and public datasets. One may hypothesize that the strong LLM (e.g. GPT-4) performance is explained by datasets being part of the LLMs’ training corpus, hence possibly being memorized. We show in Table 3 that it is unlikely, as we retain strong performance for both open-source datasets, as well as private medical datasets which require authorization processes for access and are unlikely to be part of the LLM pretraining dataset.
Remark on ICL versus fine-tuning. Our results in Table 3 and Table 4 indicate that ICL is better than fine-tuning (GReaT baseline) in the low-data regime. This highlights the difficulty of fine-tuning in this regime, where it is easy to overfit to $D_{\mathrm{train}}$. As we increase the number of samples, this baseline coupled with curation improves toward the level of CLLM  (GPT-4).
3.2 Hardness: a proxy signal to flag poor quality synthetic datasets
Figure 6: The proportion of discarded samples $D_{\mathrm{syn}}$ is a proxy for test performance. This negative linear relationship where each point is a synthetic dataset generated by GPT-4 (e.g. Adult, Covid, Compas) allows us to flag datasets that will lead to unreliable downstream performance.
Having a systematic way to assess datasets generated by LLMs like GPT-4 is important because their black-box nature provides little control on their generation quality. This contrasts conventional generators for which training loss is an exploitable signal. Hence, we ask: could we have a signal to identify a potential problematic dataset generated by GPT-4 without an exhaustive manual review? For example, GPT-4 produced low-quality synthetic data for the Adult dataset (across the different sample sizes) resulting in poor downstream performance. While curation improves it, downstream performance is still suboptimal. Addressing this question is important, since datasets are rarely created by the ML model builder in real-world ML workflows, but rather by specialist data teams or data owners (Gebru et al., 2021; Sambasivan et al., 2021; Goncalves et al., 2020). Hence, having a signal to preemptively flag a potentially suboptimal generated dataset spares investment in both storing the subpar data and/or training a model likely to underperform on real data.
$D_{\mathrm{syn}}$ should intuitively be considered imperfect if curation discards many of its samples, since the number of discarded samples measures the quality of samples with respect to the small but gold-standard $D_{\mathrm{train}}$. Hence, we investigate the relationship between test performance (AUC) and the proportion of samples discarded by the curation. Fig. 6, where each point is a synthetic dataset generated by GPT-4 (e.g. Adult, Compas), shows a strong negative linear relationship between these two quantities. This holds across the different $n$ with slopes fairly stable around $-1.4$. This relationship corroborates the poor quality of the dataset generated by GPT-4 on the Adult dataset, providing a useful proxy that $D_{\mathrm{syn}}$ is unlikely to lead to good downstream performance.
4 Discussion
We introduce CLLM, an approach for data augmentation in the ultra low-data setting. CLLM  exploits the prior knowledge of LLMs along with curation for improved downstream performance. As empirically shown, CLLM  outperforms traditional generative models—most noticeably on underrepresented subgroups, for which data augmentation is of utmost importance. CLLM  is grounded in the ICL capability of LLMs, and benefits from its simplicity. We studied GPT-3.5 and GPT-4 as backbones for CLLM. The cost of the API access pose limitations, e.g. on wide accessibility, on knowing which data was used for training the models, and on understanding the LLM’s output better. Using smaller and open LLMs could overcome these limitations, though this could come with a reduction in performance. We leave this as a promising direction for future work. Further improvements may be achieved through different tuning and prompting of the LLM, as shown in different domains (Meng et al., 2023; Liu et al., 2023). Improving LLM tuning and prompting is beyond the scope of our work, but we regard this as a promising avenue for future work.
Ethics. In this work, we evaluate CLLM  using multiple real-world datasets. The private datasets are de-identified and used in accordance with the guidance of the respective data providers. We follow recommendations to use the Azure OpenAI service when using GPT-4 and GPT-3.5 models, where via the agreement we ensure the medical data is not sent for human review or stored, hence respecting the guidelines given by the dataset providers. LLMs may make errors and may reflect or exacerbate societal biases that are present in their data (Li et al., 2023). Though the curation in CLLM improves synthetic data quality, it does not directly aim to remove biases. The quality and fairness of generated data should always be evaluated. More research into LLM bias is required before methods like CLLM should be applied to real-world sensitive settings like healthcare and finance.
Finally, increasing access to ML across regions, domains and societies is also about more than just technology. We believe broader engagement and discussion with various stakeholders is crucial to responsibly expand ML access, thereby realizing the benefits of ML in an equitable way.
Data scarcity and computational limitations are deterrents for developing ML. These challenges should inspire cutting-edge ML research (De-Arteaga et al., 2018). We believe CLLM  takes a step in this direction toward improving the use of ML in low-data settings, across society (e.g. underrepresented subgroups (Suresh and Guttag, 2021)), domains (e.g. healthcare (Alami et al., 2020; Owoyemi et al., 2020)) and regions (e.g. LMICs).
Appendix: Curated LLM: Synergy of LLMs and Data Curation for tabular augmentation in ultra low-data regimes
[…]
Appendix A Extended related work
This paper primarily engages with the work on data augmentation when we have limited data, where our primary goal is synthetic data generation to augment the dataset. Generating synthetic datasets not only helps improve downstream performance, but it is also a flexible solution as it doesn’t tie the data consumer to any particular downstream model. That said, beyond the major difference of synthetic data generation, for completeness we contrast our setting of learning from limited data with other seemingly similar settings and highlight their differences.
Contrasting learning w/ limited data vs other settings. The challenge of learning from limited data, while seemingly related to several other learning paradigms, presents distinct differences and unique intricacies that warrant dedicated study.
Transfer learning [Pan and Yang, 2009], domain adaptation [Farahani et al., 2021], and few-shot learning [Wang et al., 2020] employ additional data resources or rely on specific task-related assumptions to improve learning performance. These methods exploit large labeled data from a source domain, unlabeled data in a target domain, or leverage knowledge from related tasks respectively. For example, Levin et al. [2022] and Jin and Ucar [2023] use models trained on labeled data from a source domain, while Ruiz et al. [2023] and Margeloiu et al. [2022] leverage knowledge-graphs. This is in contrast to our setting, considered of learning with limited data, which must function with whatever scarce labeled data it has, without making any assumptions about the availability of additional data or tasks.
Active learning [Settles, 2009] and semi-supervised learning [van Engelen and Hoos, 2019, Chapelle et al., 2006] also operate under the premise of having access to plentiful unlabeled data and the capacity to interactively query labels. However, in our setting, considered learning with limited data does not inherently assume such capabilities, focusing instead on limited labeled data only.
Furthermore, active learning primarily focuses on the iterative process of selecting data samples that, when labeled, are expected to most significantly improve the model’s performance. This selection is typically based on criteria such as uncertainty sampling which focuses on epistemic uncertainty [Mussmann and Liang, 2018, Houlsby et al., 2011, Kirsch et al., 2019, Nguyen et al., 2022]. The primary objective is to minimize labeling effort while maximizing the model’s learning efficiency. Additionally, active learning would aims to label instances based on epistemic uncertainty where the model struggles to make accurate predictions, yet the samples themselves are correct. In contrast, CLLM leverage training dynamics based on aleatoric uncertainty and confidence and is designed to discard samples that might jeopardize the downstream accuracy. These samples can be considered to have inherent issues or are erroneous, such as being ”mislabeled”. To summarize, in active learning, epistemic uncertainty is used to identify data points that, if labeled, would yield the most significant insights for model training. In our approach, they serve to identify and exclude/filter data points that could potentially deteriorate the model’s performance.
Self-supervised learning [Liu et al., 2021] leverages large amounts of unlabeled data to learn useful representations for downstream tasks. However, in our setting, considered learning with limited data does not inherently assume such access to vast amounts of unlabeled data.
Data-centric AI. Ensuring high data quality is a critical but often overlooked problem in ML, where the focus is optimizing models [Sambasivan et al., 2021]. Even when it is considered, the process of assessing datasets is adhoc or artisanal [Seedat et al., 2022b]. However, the recent push of data-centric AI [Liang et al., 2022, Polyzotis and Zaharia, 2021, Zha et al., 2023] aims to develop systematic tools to curate existing datasets. Our work contributes to this nascent body of work [Seedat et al., 2023] – presenting CLLM, which, to the best of our knowledge, is the first systematic data-centric framework looking at how we can tailor synthetic datasets) rather than real datasets) to downstream task use with data curation.
Why Data Augmentation? Data augmentation is a flexible approach to address the ultra low-data regime. An alternative might be to resort to a pretrained black-box model for classification, which could be for example via in-context learning for classification [Dong et al., 2022]. However, such a solution is inadequate for several reasons, many of which would prevent real-world utility (e.g. in LMICs):
[…]
In conclusion, while large language models offer vast knowledge, for low-data settings in low-income countries, data augmentation provides a more cost-effective, controllable, and interpretable solution for building robust classifiers.
Appendix B Experimental details
We provide details on our datasets used, as well as, other experimental specifics including: generation, curation, downstream model, prompt template.
B.1 Datasets
We summarize the different datasets we use in this paper in Table 5. The datasets vary in number of samples, number of features and domain.
Table 5: Summary of the datasets used. * Denotes private/proprietary datasets.
Name $n$ samples $n$ features Domain
Adult Income [Asuncion and Newman, 2007] 30k 12 Finance
Compas [Angwin et al., 2016] 5k 13 Criminal justice
Covid-19 [Baqui et al., 2020] 7k 29 Healthcare/Medicine
CUTRACT Prostate [PCUK, 2019] 2k 12 Healthcare/Medicine
Drug [Fehrman et al., 2017] 2k 27 Healthcare/Medicine
MAGGIC [Pocock et al., 2013] 41k 29 Healthcare/Medicine
SEER Prostate [Duggan et al., 2016] 20k 12 Healthcare/Medicine
We detail the dataset splits used in Sec. 3.1. For each dataset and number of samples $n\in\{20,40,100,200\}$, we sample a training set $D_{\mathrm{train}}$ such that $|D_{\mathrm{train}}|=n$, and each target class has the same number of samples. We then split the remaining samples into two non-overlapping datasets, $D_{\mathrm{oracle}}$ and $D_{\mathrm{test}}$, which have the same cardinality. This procedure is repeated $n_{\mathrm{seed}}=10$ times, thus leading to different training and test sets. Note that the different generative models use the same $D_{\mathrm{train}}$ and $D_{\mathrm{test}}$ for a given seed.
Motivation for the choice of datasets.
  1. 1.
    Open-source: Adult, Drug and Compas are widely used open-source datasets used in the tabular data literature. Adult and Drug are both UCI datasets that have been used in many papers, while Compas is part of OpenML Vanschoren et al. [2013]. Our reason for selecting them is that, despite them being open-source, they are highly reflective of domains in which we might be unable to collect many samples — hence in reality would often be in an ultra-low data regime.
  2. 2.
    Private datasets: We wanted to disentangle the possible role of memorization in the strong performance of the LLM. To ensure the datasets are not in the LLMs training corpus, we selected 4 private medical datasets that need an authorization process to access. Hence, these datasets would not be part of the LLMs training corpus given their proprietary nature and hence would be unseen to the LLM. While the private and unseen aspect was the main motivation, we also wish to highlight that these are real-world medical datasets. Consequently, this allows us to test a highly realistic problem setting.
B.2 Data generation.
GPT-4 and GPT-3.5 We access GPT-4 [OpenAI, 2023] and GPT-3.5-Turbo [Brown et al., 2020] through the API. We use a temperature of $0.9$.
GReaT. GReaT Borisov et al. [2023] is a generative model which fine-tunes an LLM based on a training set. We use the implementation provided by authors.
Generative model based approaches. For the other baselines used in 3.1, we use the library SynthCity [Qian et al., 2023], using the defaults. We detail each next.
  • TVAE: this is a conditional Variational Auto Encoder (VAE) for tabular data and is based on Xu et al. [2019]
  • CTGAN: A conditional generative adversarial network which can handle tabular data and is based on Xu et al. [2019]
  • NFLOW: Normalizing Flows are generative models which produce tractable distributions where both sampling and density evaluation can be efficient and exact.
  • TabDDPM: A diffusion model that can be universally applied to any tabular dataset, handles any type of feature and is based on Kotelnikov et al. [2022]
Traditional Data Augmentation. We use SMOTE [Chawla et al., 2002] which augments data by considering nearest neighbors and performing linear interpolations. We use the implementation provided by Lemaître et al. [2017], and set the number of neighbors $k$ to $5$.
B.3 Data curation
Learning dynamics computation
We train an XGBoost with $100$ estimators on $D_{\mathrm{train}}$. We then compute predictive confidence and aleatoric uncertainty for the samples in $D_{\mathrm{syn}}$. The motivation for the choice of an XGBoost backbone is that we cannot expect good performance by choosing “any” curation model, but rather we require a curation model with enough capacity and generalization properties — where boosting methods like XGBoost used in our work have shown to achieve best performance on tabular data. This leads to our guideline for the curation step: the model used for curation should be at least as flexible as the model that the practitioner intends to use for the downstream task.
Learning dynamics thresholds
Recall that CLLM  has two thresholds $\tau_{\mathrm{conf}}$ and $\tau_{\mathrm{al}}$ on the predictive confidence and aleatoric uncertainty respectively, as defined in 2.2. We set $\tau_{\mathrm{conf}}=0.2$, in order to select high confidence samples. We adopt an adaptive threshold for $\tau_{\mathrm{al}}$ based on the dataset, such that $\tau_{\mathrm{al}}=0.75\cdot(\max(v_{al}(D_{\mathrm{syn}}))-\min(v_{al}(D_{\mathrm{syn}})))$. Note that by definition $v_{al}(D_{\mathrm{syn}})$ is bounded between 00 and $0.25$.
Example of learning dynamics
We include examples of learning dynamics computed for $20$ samples in Fig. 7.
Figure 7: Learning dynamics computed for $20$ samples
B.4 Downstream task
We compute downstream performance in Sec. 3.1 using four different downstream models: XGBoost, Random Forest, Decision tree, and Logistic Regression.
B.5 Prompt example
We include the template of the prompts used throughout the paper. We show how we include (1) in-context examples (demonstrations), (2) contextual information including dataset background and feature information and (3) the instruction.
1    System role: ’You are a tabular synthetic data generation model.’
2
3    You are a synthetic data generator.
4    Your goal is to produce data which mirrors \
5    the given examples in causal structure and feature and label distributions \
6    but also produce as diverse samples as possible.
7
8    I will give you real examples first.
9
10    Context: Leverage your medical knowledge about covid and Brazil to generate 1000 realistic but diverse samples.
11
12    example data: {data}
13
14    The output should be a markdown code snippet formatted in the following schema:
15
16    "Sex_male": string  // feature column
17    "Age": string  // feature column
18    "Age_40": string  // feature column
19    "Age_40_50": string  // feature column
20    "Age_50_60": string  // feature column
21    "Age_60_70": string  // feature column
22    "Age_70": string  // feature column
23    "Fever": string  // feature column
24    "Cough": string  // feature column
25    "Sore_throat": string  // feature column
26    "Shortness_of_breath": string  // feature column
27    "Respiratory_discomfort": string  // feature column
28    "SPO2": string  // feature column
29    "Dihareea": string  // feature column
30    "Vomitting": string  // feature column
31    "Cardiovascular": string  // feature column
32    "Asthma": string  // feature column
33    "Diabetis": string  // feature column
34    "Pulmonary": string  // feature column
35    "Immunosuppresion": string  // feature column
36    "Obesity": string  // feature column
37    "Liver": string  // feature column
38    "Neurologic": string  // feature column
39    "Renal": string  // feature column
40    "Branca": string  // feature column
41    "Preta": string  // feature column
42    "Amarela": string  // feature column
43    "Parda": string  // feature column
44    "Indigena": string  // feature column
45    "is_dead": string  // label if patient dead or not, is_dead
46
47    DO NOT COPY THE EXAMPLES but generate realistic but new and diverse samples which have the correct label conditioned on the features.
Listing 1: Template of the prompt
Appendix C Additional results
C.1 Detailed results for Section 3.1
We report additional results for Sec. 3.1, showing the AUC for each downstream model (XGBoost, Random forest, Logistic regression, Decision tree). As we can see, the conclusion that curation helps improve downstream performance holds for each of these various downstream models, as is indicated by the green arrows in Tables 6, 7, 8, 9.
Table 6: AUC for the RF model on $D_{\mathrm{test}}$ where curation improves performance for all methods across all sample sizes $n$, as indicated by $\uparrow$.
Real data CLLM (OURS) Baselines
GPT-4 GPT-3.5 CTGAN TabDDPM GReaT NFLOW SMOTE TVAE
Dataset $D_{\mathrm{oracle}}$ $D_{\mathrm{train}}$ Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur.
covid (n=20) 76.11 72.52 75.67 75.59 72.37 73.32 $\uparrow$ 61.66 65.24 $\uparrow$ 70.22 70.13 57.78 68.24 $\uparrow$ 65.54 70.72 $\uparrow$ 71.40 71.42 $\uparrow$ 63.90 67.69 $\uparrow$
cutract (n=20) 74.16 73.34 73.45 74.24 normal-$\uparrow$ 72.22 73.51 $\uparrow$ 66.24 70.45 $\uparrow$ 69.73 70.49 $\uparrow$ 52.61 68.11 $\uparrow$ 66.63 72.39 $\uparrow$ 71.42 72.34 $\uparrow$ 71.18 71.89 $\uparrow$
maggic (n=20) 71.28 59.18 62.92 63.97 normal-$\uparrow$ 60.10 61.25 $\uparrow$ 53.21 55.14 $\uparrow$ 55.61 56.50 $\uparrow$ 50.92 57.23 $\uparrow$ 55.91 59.07 $\uparrow$ 58.20 58.55 $\uparrow$ 55.43 57.71 $\uparrow$
seer (n=20) 90.09 85.22 86.30 86.82 normal-$\uparrow$ 85.86 85.73 78.11 80.52 $\uparrow$ 82.61 82.74 $\uparrow$ 47.57 75.70 $\uparrow$ 77.35 81.45 $\uparrow$ 82.09 83.00 $\uparrow$ 77.63 81.16 $\uparrow$
compas (n=20) 66.79 64.47 68.40 67.85 61.93 64.29 $\uparrow$ 56.87 63.21 $\uparrow$ 58.56 61.26 $\uparrow$ 52.23 63.30 $\uparrow$ 59.60 64.79 $\uparrow$ 61.05 62.18 $\uparrow$ 59.31 63.37 $\uparrow$
adult (n=20) 85.83 83.31 51.35 73.45 $\uparrow$ 49.17 74.27 $\uparrow$ 75.63 78.99 $\uparrow$ 77.27 77.64 $\uparrow$ 69.35 78.79 $\uparrow$ 68.32 78.27 $\uparrow$ 77.75 78.34 $\uparrow$ 76.13 79.02 $\uparrow$
drug (n=20) 83.12 77.07 79.23 78.90 77.11 76.95 72.81 77.24 $\uparrow$ 72.00 74.03 $\uparrow$ 63.34 73.29 $\uparrow$ 65.47 70.19 $\uparrow$ 75.48 75.36 72.22 74.32 $\uparrow$
covid (n=40) 76.71 75.21 75.63 75.70 normal-$\uparrow$ 72.27 73.59 $\uparrow$ 66.64 70.75 $\uparrow$ 75.56 75.39 56.91 70.88 $\uparrow$ 70.47 73.15 $\uparrow$ 73.59 73.63 $\uparrow$ 64.27 69.71 $\uparrow$
cutract (n=40) 74.96 72.31 71.67 73.20 normal-$\uparrow$ 69.92 71.01 $\uparrow$ 64.67 69.21 $\uparrow$ 69.16 69.67 $\uparrow$ 53.41 68.93 $\uparrow$ 61.23 69.21 $\uparrow$ 70.82 71.21 $\uparrow$ 59.99 67.56 $\uparrow$
maggic (n=40) 71.32 60.91 61.80 63.11 normal-$\uparrow$ 59.38 61.50 $\uparrow$ 56.49 58.17 $\uparrow$ 56.50 58.21 $\uparrow$ 48.43 57.84 $\uparrow$ 55.47 59.94 $\uparrow$ 59.82 60.11 $\uparrow$ 56.76 58.92 $\uparrow$
seer (n=40) 90.06 86.96 86.01 86.29 $\uparrow$ 86.73 87.09 normal-$\uparrow$ 83.46 86.62 $\uparrow$ 85.86 85.19 54.71 83.22 $\uparrow$ 83.23 86.81 $\uparrow$ 83.76 84.82 $\uparrow$ 80.34 85.82 $\uparrow$
compas (n=40) 66.66 62.73 68.06 68.13 normal-$\uparrow$ 61.17 62.27 $\uparrow$ 56.05 60.92 $\uparrow$ 59.76 61.03 $\uparrow$ 57.82 64.25 $\uparrow$ 58.89 63.79 $\uparrow$ 61.20 61.25 $\uparrow$ 55.89 60.54 $\uparrow$
adult (n=40) 86.04 83.61 50.26 75.55 $\uparrow$ 46.68 76.63 $\uparrow$ 75.13 81.56 $\uparrow$ 76.10 80.40 $\uparrow$ 68.04 81.08 $\uparrow$ 73.29 81.46 $\uparrow$ 80.35 81.10 $\uparrow$ 75.81 82.11 $\uparrow$
drug (n=40) 82.76 78.81 78.35 79.54 normal-$\uparrow$ 77.33 77.94 $\uparrow$ 74.68 76.64 $\uparrow$ 74.50 77.18 $\uparrow$ 70.25 $7$uparrow 65.86 74.55 $\uparrow$ 74.84 75.26 $\uparrow$ 71.29 75.98 $\uparrow$
covid (n=100) 76.21 75.78 75.86 76.33 normal-$\uparrow$ 73.02 74.40 $\uparrow$ 72.00 74.74 $\uparrow$ 74.64 75.74 $\uparrow$ 65.76 74.69 $\uparrow$ 67.05 75.47 $\uparrow$ 74.00 74.09 $\uparrow$ 72.11 74.44 $\uparrow$
cutract (n=100) 74.59 73.93 72.92 74.73 $\uparrow$ 72.51 73.93 $\uparrow$ 70.18 74.26 $\uparrow$ 71.71 73.69 $\uparrow$ 55.48 70.79 $\uparrow$ 69.79 74.75 normal-$\uparrow$ 71.18 72.47 $\uparrow$ 68.47 73.78 $\uparrow$
maggic (n=100) 71.28 63.06 60.99 63.36 normal-$\uparrow$ 57.97 60.86 $\uparrow$ 58.98 60.76 $\uparrow$ 58.88 60.13 $\uparrow$ 49.66 59.57 $\uparrow$ 57.82 62.24 $\uparrow$ 61.55 61.90 $\uparrow$ 57.46 60.55 $\uparrow$
seer (n=100) 89.94 87.53 86.33 87.31 $\uparrow$ 86.40 87.06 $\uparrow$ 84.59 87.27 $\uparrow$ 85.89 87.00 $\uparrow$ 70.32 85.97 $\uparrow$ 84.30 87.52 $\uparrow$ 85.12 85.82 $\uparrow$ 81.92 86.41 $\uparrow$
compas (n=100) 66.38 63.18 68.44 68.67 normal-$\uparrow$ 59.41 62.00 $\uparrow$ 60.28 64.34 $\uparrow$ 60.17 62.63 $\uparrow$ 59.32 63.98 $\uparrow$ 60.10 65.19 $\uparrow$ 61.34 61.37 $\uparrow$ 59.58 63.37 $\uparrow$
adult (n=100) 85.99 84.38 46.37 76.29 $\uparrow$ 47.30 75.74 $\uparrow$ 77.45 82.63 $\uparrow$ 80.88 81.69 $\uparrow$ 79.23 83.33 $\uparrow$ 73.52 83.38 $\uparrow$ 82.54 83.05 $\uparrow$ 76.60 82.28 $\uparrow$
drug (n=100) 83.18 80.31 79.29 79.75 $\uparrow$ 74.55 76.46 $\uparrow$ 74.54 78.01 $\uparrow$ 78.34 79.88 $\uparrow$ 73.19 79.72 $\uparrow$ 67.01 76.46 $\uparrow$ 76.92 77.57 $\uparrow$ 74.53 78.82 $\uparrow$
covid (n=200) 76.31 76.08 75.23 75.84 $\uparrow$ 72.62 74.99 $\uparrow$ 74.04 76.25 $\uparrow$ 75.28 76.82 normal-$\uparrow$ 67.08 75.62 $\uparrow$ 67.73 75.67 $\uparrow$ 74.82 75.27 $\uparrow$ 70.64 75.03 $\uparrow$
cutract (n=200) 75.04 74.25 73.48 75.28 $\uparrow$ 72.01 74.37 $\uparrow$ 71.99 74.93 $\uparrow$ 74.33 76.19 normal-$\uparrow$ 68.05 74.68 $\uparrow$ 71.47 75.67 $\uparrow$ 72.44 72.85 $\uparrow$ 68.55 74.43 $\uparrow$
maggic (n=200) 71.07 64.77 61.56 63.78 $\uparrow$ 59.76 62.53 $\uparrow$ 60.85 63.57 $\uparrow$ 58.23 58.85 $\uparrow$ 50.84 61.49 $\uparrow$ 57.52 63.00 $\uparrow$ 63.72 64.01 $\uparrow$ 58.92 62.08 $\uparrow$
seer (n=200) 89.99 88.13 86.96 87.72 $\uparrow$ 85.28 86.70 $\uparrow$ 85.05 87.83 $\uparrow$ 87.14 88.80 normal-$\uparrow$ 84.86 88.14 $\uparrow$ 80.80 87.46 $\uparrow$ 86.57 87.11 $\uparrow$ 82.30 87.59 $\uparrow$
compas (n=200) 66.22 63.61 67.68 68.80 normal-$\uparrow$ 60.67 63.43 $\uparrow$ 61.20 65.11 $\uparrow$ 60.88 63.87 $\uparrow$ 60.10 64.91 $\uparrow$ 56.94 64.76 $\uparrow$ 61.57 62.76 $\uparrow$ 59.95 64.14 $\uparrow$
adult (n=200) 86.01 85.16 40.18 77.61 $\uparrow$ 46.74 78.71 $\uparrow$ 80.92 84.26 $\uparrow$ 84.50 85.44 normal-$\uparrow$ 82.73 85.05 $\uparrow$ 77.48 84.72 $\uparrow$ 84.26 84.33 $\uparrow$ 78.12 83.94 $\uparrow$
drug (n=200) 82.14 81.47 78.70 78.98 $\uparrow$ 75.87 77.59 $\uparrow$ 77.19 80.03 $\uparrow$ 73.25 75.75 $\uparrow$ 78.50 81.13 $\uparrow$ 67.83 78.81 $\uparrow$ 79.50 80.16 $\uparrow$ 74.81 79.45 $\uparrow$
Table 7: AUC for the XGB model on $D_{\mathrm{test}}$ where curation improves performance for all methods across all sample sizes $n$, as indicated by $\uparrow$.
Real data CLLM (OURS) Baselines
GPT-4 GPT-3.5 CTGAN TabDDPM GReaT NFLOW SMOTE TVAE
Dataset $D_{\mathrm{oracle}}$ $D_{\mathrm{train}}$ Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur.
covid (n=20) 77.59 69.70 75.59 75.49 71.27 72.25 $\uparrow$ 61.40 65.96 $\uparrow$ 69.39 68.21 55.99 67.06 $\uparrow$ 62.28 69.67 $\uparrow$ 68.49 68.43 62.49 67.16 $\uparrow$
cutract (n=20) 73.14 71.04 73.26 74.30 normal-$\uparrow$ 71.62 73.04 $\uparrow$ 65.13 69.67 $\uparrow$ 68.46 68.87 $\uparrow$ 51.99 68.31 $\uparrow$ 62.93 70.87 $\uparrow$ 68.28 70.72 $\uparrow$ 69.96 71.98 $\uparrow$
maggic (n=20) 71.38 57.98 61.54 62.81 normal-$\uparrow$ 58.76 60.01 $\uparrow$ 53.62 55.44 $\uparrow$ 55.97 57.06 $\uparrow$ 50.14 56.96 $\uparrow$ 54.91 58.22 $\uparrow$ 56.79 57.15 $\uparrow$ 55.41 57.72 $\uparrow$
seer (n=20) 91.80 80.41 87.00 86.56 85.67 85.42 76.84 80.00 $\uparrow$ 81.79 80.21 45.06 75.39 $\uparrow$ 76.25 81.45 $\uparrow$ 78.57 79.78 $\uparrow$ 75.69 81.01 $\uparrow$
compas (n=20) 69.17 65.49 69.71 68.79 62.55 64.68 $\uparrow$ 56.34 63.92 $\uparrow$ 58.19 61.42 $\uparrow$ 51.73 63.51 $\uparrow$ 59.01 64.84 $\uparrow$ 62.35 63.46 $\uparrow$ 59.18 63.62 $\uparrow$
adult (n=20) 88.64 79.21 52.12 74.28 $\uparrow$ 47.70 72.91 $\uparrow$ 71.70 78.26 $\uparrow$ 76.50 77.22 $\uparrow$ 67.79 78.67 $\uparrow$ 66.79 $7$uparrow 77.11 77.97 $\uparrow$ 75.46 79.35 normal-$\uparrow$
drug (n=20) 81.06 69.26 78.17 77.28 75.23 75.71 $\uparrow$ 70.60 74.31 $\uparrow$ 70.29 71.25 $\uparrow$ 57.18 69.67 $\uparrow$ 62.96 69.12 $\uparrow$ 70.84 72.30 $\uparrow$ 69.78 72.45 $\uparrow$
covid (n=40) 77.93 71.57 75.23 75.36 normal-$\uparrow$ 71.55 73.37 $\uparrow$ 66.17 70.85 $\uparrow$ 72.21 72.68 $\uparrow$ 56.03 70.80 $\uparrow$ 67.39 71.82 $\uparrow$ 70.02 69.96 61.19 69.77 $\uparrow$
cutract (n=40) 73.83 69.50 71.26 73.57 normal-$\uparrow$ 69.66 71.34 $\uparrow$ 62.52 68.69 $\uparrow$ 64.50 68.05 $\uparrow$ 51.23 68.23 $\uparrow$ 60.74 68.91 $\uparrow$ 69.42 69.19 59.34 67.80 $\uparrow$
maggic (n=40) 71.45 58.76 60.27 61.88 normal-$\uparrow$ 57.87 60.29 $\uparrow$ 55.87 58.47 $\uparrow$ 55.38 58.03 $\uparrow$ 48.50 57.02 $\uparrow$ 54.53 59.01 $\uparrow$ 58.12 58.29 $\uparrow$ 56.19 58.33 $\uparrow$
seer (n=40) 91.79 84.51 86.63 87.40 normal-$\uparrow$ 86.27 86.80 $\uparrow$ 82.48 86.55 $\uparrow$ 81.62 81.58 52.80 83.03 $\uparrow$ 80.04 85.84 $\uparrow$ 82.53 83.62 $\uparrow$ 79.35 85.81 $\uparrow$
compas (n=40) 69.26 62.90 69.24 68.90 61.74 63.21 $\uparrow$ 56.63 62.55 $\uparrow$ 59.07 59.38 $\uparrow$ 58.01 64.95 $\uparrow$ 58.44 64.44 $\uparrow$ 61.81 61.89 $\uparrow$ 56.60 62.26 $\uparrow$
adult (n=40) 88.88 82.16 48.51 76.59 $\uparrow$ 45.92 76.92 $\uparrow$ 73.38 82.53 $\uparrow$ 74.03 79.84 $\uparrow$ 66.60 80.98 $\uparrow$ 69.47 81.56 $\uparrow$ 80.01 80.51 $\uparrow$ 74.20 82.53 normal-$\uparrow$
drug (n=40) 80.63 72.72 77.03 78.85 normal-$\uparrow$ 76.37 76.48 $\uparrow$ 70.99 74.48 $\uparrow$ 74.65 76.53 $\uparrow$ 64.52 75.09 $\uparrow$ 62.58 73.14 $\uparrow$ 71.80 72.15 $\uparrow$ 67.61 74.04 $\uparrow$
covid (n=100) 77.70 73.16 75.55 76.56 normal-$\uparrow$ 72.69 74.36 $\uparrow$ 70.54 73.99 $\uparrow$ 72.99 74.50 $\uparrow$ 63.40 73.98 $\uparrow$ 63.50 73.79 $\uparrow$ 72.64 73.35 $\uparrow$ 70.45 73.76 $\uparrow$
cutract (n=100) 73.04 70.49 72.29 74.01 $\uparrow$ 71.51 73.48 $\uparrow$ 68.64 73.63 $\uparrow$ 71.32 74.02 normal-$\uparrow$ 54.74 69.69 $\uparrow$ 66.97 72.98 $\uparrow$ 67.71 69.42 $\uparrow$ 66.44 72.32 $\uparrow$
maggic (n=100) 71.42 60.55 60.04 62.57 normal-$\uparrow$ 56.98 60.49 $\uparrow$ 58.21 60.42 $\uparrow$ 58.11 59.42 $\uparrow$ 49.71 58.77 $\uparrow$ 55.36 60.58 $\uparrow$ 60.02 60.15 $\uparrow$ 56.47 60.28 $\uparrow$
seer (n=100) 91.81 85.52 86.81 87.64 normal-$\uparrow$ 86.39 87.21 $\uparrow$ 83.74 87.29 $\uparrow$ 84.66 86.85 $\uparrow$ 68.20 85.49 $\uparrow$ 82.37 86.80 $\uparrow$ 83.18 83.87 $\uparrow$ 80.45 86.49 $\uparrow$
compas (n=100) 68.88 62.71 69.37 69.37 59.83 62.80 $\uparrow$ 60.56 64.90 $\uparrow$ 58.20 62.00 $\uparrow$ 58.51 64.60 $\uparrow$ 59.76 64.68 $\uparrow$ 61.33 62.03 $\uparrow$ 60.37 63.42 $\uparrow$
adult (n=100) 88.71 82.78 44.60 76.75 $\uparrow$ 46.34 76.83 $\uparrow$ 74.97 82.89 $\uparrow$ 77.38 81.42 $\uparrow$ 78.84 83.96 normal-$\uparrow$ 68.31 82.87 $\uparrow$ 82.58 83.24 $\uparrow$ 76.15 82.47 $\uparrow$
drug (n=100) 81.44 75.49 79.12 79.81 normal-$\uparrow$ 72.52 74.78 $\uparrow$ 68.89 75.91 $\uparrow$ 76.88 78.33 $\uparrow$ 66.38 77.02 $\uparrow$ 62.27 75.13 $\uparrow$ 74.40 74.49 $\uparrow$ 71.02 76.85 $\uparrow$
covid (n=200) 77.56 73.06 75.12 76.64 normal-$\uparrow$ 71.74 74.66 $\uparrow$ 72.57 75.50 $\uparrow$ 73.56 75.88 $\uparrow$ 66.02 74.51 $\uparrow$ 63.29 74.62 $\uparrow$ 73.34 73.89 $\uparrow$ 68.14 74.21 $\uparrow$
cutract (n=200) 73.68 71.68 73.43 75.14 $\uparrow$ 71.76 74.31 $\uparrow$ 69.35 73.80 $\uparrow$ 72.76 75.33 normal-$\uparrow$ 65.83 73.43 $\uparrow$ 68.05 74.27 $\uparrow$ 70.06 71.35 $\uparrow$ 66.65 73.34 $\uparrow$
maggic (n=200) 71.41 62.40 59.46 63.15 normal-$\uparrow$ 58.05 61.68 $\uparrow$ 59.04 62.58 $\uparrow$ 57.63 59.23 $\uparrow$ 50.63 60.09 $\uparrow$ 55.55 61.36 $\uparrow$ 61.73 61.88 $\uparrow$ 57.09 61.15 $\uparrow$
seer (n=200) 91.83 85.87 86.75 88.08 $\uparrow$ 85.74 86.97 $\uparrow$ 83.97 87.64 $\uparrow$ 86.21 88.46 normal-$\uparrow$ 83.21 87.65 $\uparrow$ 76.91 87.16 $\uparrow$ 84.70 85.59 $\uparrow$ 81.45 87.46 $\uparrow$
compas (n=200) 69.04 63.14 68.80 69.54 normal-$\uparrow$ 61.62 64.59 $\uparrow$ 60.65 65.67 $\uparrow$ 60.97 65.12 $\uparrow$ 60.66 65.67 $\uparrow$ 56.49 64.34 $\uparrow$ 61.68 62.93 $\uparrow$ 60.32 64.18 $\uparrow$
adult (n=200) 88.86 84.12 40.38 77.39 $\uparrow$ 47.40 79.40 $\uparrow$ 80.00 84.61 $\uparrow$ 83.73 85.73 $\uparrow$ 81.92 85.76 normal-$\uparrow$ 74.90 85.13 $\uparrow$ 84.78 84.45 75.54 84.39 $\uparrow$
drug (n=200) 80.13 78.41 78.25 79.09 $\uparrow$ 73.31 76.76 $\uparrow$ 72.82 79.00 $\uparrow$ 72.62 74.89 $\uparrow$ 74.30 79.75 normal-$\uparrow$ 63.91 77.60 $\uparrow$ 77.64 77.92 $\uparrow$ 70.16 77.89 $\uparrow$
Table 8: AUC for the DT model on $D_{\mathrm{test}}$ where curation improve performance for all methods across all sample sizes $n$, as indicated by $\uparrow$.
Real data CLLM (OURS) Baselines
GPT-4 GPT-3.5 CTGAN TabDDPM GReaT NFLOW SMOTE TVAE
Dataset $D_{\mathrm{oracle}}$ $D_{\mathrm{train}}$ Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur.
covid (n=20) 63.46 61.92 67.52 67.67 normal-$\uparrow$ 61.07 64.89 $\uparrow$ 52.94 56.62 $\uparrow$ 58.95 59.19 $\uparrow$ 51.83 58.73 $\uparrow$ 54.37 61.53 $\uparrow$ 59.56 59.01 56.46 59.49 $\uparrow$
cutract (n=20) 62.49 64.62 62.40 66.03 normal-$\uparrow$ 62.10 65.15 $\uparrow$ 55.67 61.73 $\uparrow$ 57.91 58.97 $\uparrow$ 51.89 62.76 $\uparrow$ 57.12 64.85 $\uparrow$ 62.85 62.38 60.25 62.34 $\uparrow$
maggic (n=20) 56.53 54.80 54.67 55.03 normal-$\uparrow$ 52.54 53.54 $\uparrow$ 50.90 52.08 $\uparrow$ 52.39 52.70 $\uparrow$ 49.70 53.08 $\uparrow$ 51.52 53.23 $\uparrow$ 52.55 52.44 50.91 51.52 $\uparrow$
seer (n=20) 78.86 71.09 76.17 77.17 normal-$\uparrow$ 73.92 75.58 $\uparrow$ 64.18 70.69 $\uparrow$ 74.78 74.72 46.60 68.79 $\uparrow$ 65.77 72.25 $\uparrow$ 70.54 71.19 $\uparrow$ 64.47 70.11 $\uparrow$
compas (n=20) 61.05 58.73 62.07 63.91 normal-$\uparrow$ 57.17 60.85 $\uparrow$ 53.83 58.83 $\uparrow$ 51.77 56.11 $\uparrow$ 52.39 60.96 $\uparrow$ 54.64 60.27 $\uparrow$ 58.16 57.57 53.52 58.97 $\uparrow$
adult (n=20) 73.84 68.03 52.80 67.10 $\uparrow$ 49.15 66.52 $\uparrow$ 61.57 68.93 $\uparrow$ 70.46 71.19 $\uparrow$ 62.04 71.87 normal-$\uparrow$ 59.07 69.16 $\uparrow$ 65.52 63.33 64.40 70.41 $\uparrow$
drug (n=20) 65.66 63.49 64.40 66.87 normal-$\uparrow$ 62.69 64.21 $\uparrow$ 56.76 64.18 $\uparrow$ 61.04 62.48 $\uparrow$ 52.22 62.62 $\uparrow$ 53.50 61.30 $\uparrow$ 62.34 60.69 55.00 58.89 $\uparrow$
covid (n=40) 64.94 65.10 66.55 67.85 normal-$\uparrow$ 62.54 64.64 $\uparrow$ 57.16 63.41 $\uparrow$ 63.15 60.93 51.66 61.81 $\uparrow$ 57.04 63.13 $\uparrow$ 62.19 60.49 54.51 60.32 $\uparrow$
cutract (n=40) 62.32 63.01 62.49 65.29 normal-$\uparrow$ 61.61 63.01 $\uparrow$ 56.40 62.63 $\uparrow$ 58.18 58.97 $\uparrow$ 51.41 62.50 $\uparrow$ 55.64 62.24 $\uparrow$ 60.93 61.29 $\uparrow$ 51.94 60.24 $\uparrow$
maggic (n=40) 56.85 53.74 53.20 54.71 normal-$\uparrow$ 52.33 52.97 $\uparrow$ 51.85 53.85 $\uparrow$ 52.48 53.45 $\uparrow$ 49.61 52.87 $\uparrow$ 51.24 54.45 $\uparrow$ 53.69 53.37 50.60 52.66 $\uparrow$
seer (n=40) 78.84 72.40 76.02 77.13 $\uparrow$ 75.18 78.26 normal-$\uparrow$ 68.31 73.89 $\uparrow$ 77.94 73.24 51.45 72.32 $\uparrow$ 67.78 75.95 $\uparrow$ 70.32 72.81 $\uparrow$ 68.79 74.48 $\uparrow$
compas (n=40) 61.12 58.23 62.00 63.43 normal-$\uparrow$ 56.73 59.19 $\uparrow$ 52.23 56.84 $\uparrow$ 56.67 57.55 $\uparrow$ 54.52 60.09 $\uparrow$ 54.47 58.92 $\uparrow$ 56.59 56.47 52.62 57.20 $\uparrow$
adult (n=40) 74.13 69.75 49.47 68.90 $\uparrow$ 47.85 66.79 $\uparrow$ 60.55 70.14 $\uparrow$ 64.06 70.56 $\uparrow$ 58.77 71.63 $\uparrow$ 60.26 71.50 $\uparrow$ 69.43 70.58 $\uparrow$ 59.76 72.41 normal-$\uparrow$
drug (n=40) 66.19 64.18 62.86 65.90 normal-$\uparrow$ 60.09 63.43 $\uparrow$ 60.84 65.40 $\uparrow$ 63.05 64.90 $\uparrow$ 54.80 65.25 $\uparrow$ 54.68 62.40 $\uparrow$ 61.21 61.98 $\uparrow$ 53.77 60.23 $\uparrow$
covid (n=100) 63.57 63.17 67.14 68.60 normal-$\uparrow$ 62.40 66.02 $\uparrow$ 61.34 64.63 $\uparrow$ 63.33 66.42 $\uparrow$ 55.86 64.53 $\uparrow$ 55.44 64.91 $\uparrow$ 61.96 63.40 $\uparrow$ 59.75 64.47 $\uparrow$
cutract (n=100) 62.88 62.69 60.96 65.60 $\uparrow$ 62.12 65.95 $\uparrow$ 59.06 66.20 normal-$\uparrow$ 63.38 65.72 $\uparrow$ 52.62 61.73 $\uparrow$ 58.11 65.08 $\uparrow$ 60.15 61.34 $\uparrow$ 58.82 63.76 $\uparrow$
maggic (n=100) 56.67 54.83 53.58 55.70 normal-$\uparrow$ 50.72 53.98 $\uparrow$ 52.29 55.06 $\uparrow$ 53.93 54.61 $\uparrow$ 49.71 53.34 $\uparrow$ 51.82 54.53 $\uparrow$ 54.11 54.34 $\uparrow$ 51.31 53.33 $\uparrow$
seer (n=100) 78.48 73.83 76.50 77.84 normal-$\uparrow$ 74.81 77.80 $\uparrow$ 69.61 76.91 $\uparrow$ 74.07 77.21 $\uparrow$ 59.62 75.40 $\uparrow$ 69.51 76.54 $\uparrow$ 71.04 72.63 $\uparrow$ 66.35 76.44 $\uparrow$
compas (n=100) 61.17 56.61 63.04 63.39 normal-$\uparrow$ 55.57 57.40 $\uparrow$ 55.07 58.95 $\uparrow$ 53.38 55.69 $\uparrow$ 54.61 59.31 $\uparrow$ 54.38 57.92 $\uparrow$ 56.69 57.01 $\uparrow$ 54.34 56.72 $\uparrow$
adult (n=100) 74.22 72.19 46.68 66.58 $\uparrow$ 47.94 67.37 $\uparrow$ 62.42 71.78 $\uparrow$ 68.06 70.77 $\uparrow$ 66.29 74.36 normal-$\uparrow$ 58.28 72.14 $\uparrow$ 71.51 72.68 $\uparrow$ 62.11 72.18 $\uparrow$
drug (n=100) 65.98 64.20 66.18 67.41 normal-$\uparrow$ 60.77 61.21 $\uparrow$ 57.67 63.41 $\uparrow$ 63.26 64.55 $\uparrow$ 54.89 65.72 $\uparrow$ 52.06 63.38 $\uparrow$ 61.21 62.30 $\uparrow$ 55.40 63.32 $\uparrow$
covid (n=200) 64.28 63.45 67.19 68.74 normal-$\uparrow$ 63.25 65.73 $\uparrow$ 61.58 66.27 $\uparrow$ 63.55 66.98 $\uparrow$ 57.20 65.54 $\uparrow$ 55.08 63.77 $\uparrow$ 63.60 63.40 60.01 64.27 $\uparrow$
cutract (n=200) 62.77 63.53 63.36 65.33 $\uparrow$ 63.01 65.47 $\uparrow$ 60.51 65.01 $\uparrow$ 62.55 66.94 normal-$\uparrow$ 57.40 65.14 $\uparrow$ 59.13 65.20 $\uparrow$ 61.66 62.40 $\uparrow$ 58.29 64.85 $\uparrow$
maggic (n=200) 56.72 55.06 52.76 54.41 $\uparrow$ 52.41 54.34 $\uparrow$ 52.47 55.21 normal-$\uparrow$ 52.40 53.39 $\uparrow$ 50.43 53.66 $\uparrow$ 51.58 54.82 $\uparrow$ 54.64 54.36 51.78 54.22 $\uparrow$
seer (n=200) 78.58 74.51 75.20 77.26 $\uparrow$ 75.33 77.41 $\uparrow$ 70.96 77.08 $\uparrow$ 74.51 79.20 normal-$\uparrow$ 71.44 78.66 $\uparrow$ 65.84 76.60 $\uparrow$ 73.11 74.77 $\uparrow$ 66.37 76.36 $\uparrow$
compas (n=200) 60.82 57.20 61.09 62.89 normal-$\uparrow$ 56.15 59.02 $\uparrow$ 54.20 57.78 $\uparrow$ 55.72 58.68 $\uparrow$ 55.83 59.81 $\uparrow$ 52.13 57.25 $\uparrow$ 56.55 57.34 $\uparrow$ 54.07 58.66 $\uparrow$
adult (n=200) 73.61 72.26 44.41 67.99 $\uparrow$ 48.46 67.24 $\uparrow$ 66.56 74.12 $\uparrow$ 71.61 74.89 $\uparrow$ 69.91 76.52 normal-$\uparrow$ 60.70 74.05 $\uparrow$ 73.11 74.08 $\uparrow$ 63.10 73.82 $\uparrow$
drug (n=200) 65.85 67.04 66.29 66.57 $\uparrow$ 61.88 64.34 $\uparrow$ 59.10 66.06 $\uparrow$ 63.59 63.40 59.11 69.04 normal-$\uparrow$ 53.78 64.55 $\uparrow$ 64.56 65.35 $\uparrow$ 56.82 63.02 $\uparrow$
Table 9: AUC for the LR model on $D_{\mathrm{test}}$ where curation improves performance for all methods across all sample sizes $n$, as indicated by $\uparrow$.
Real data CLLM (OURS) Baselines
GPT-4 GPT-3.5 CTGAN TabDDPM GReaT NFLOW SMOTE TVAE
Dataset $D_{\mathrm{oracle}}$ $D_{\mathrm{train}}$ Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur.
covid (n=20) 80.47 69.85 76.35 76.73 normal-$\uparrow$ 74.69 75.18 $\uparrow$ 59.99 66.87 $\uparrow$ 68.78 69.84 $\uparrow$ 63.94 71.81 $\uparrow$ 69.28 72.32 $\uparrow$ 68.36 68.44 $\uparrow$ 63.92 70.12 $\uparrow$
cutract (n=20) 79.12 71.47 75.50 75.41 73.93 74.47 $\uparrow$ 69.01 70.06 $\uparrow$ 68.11 68.02 53.04 68.90 $\uparrow$ 71.09 73.55 $\uparrow$ 71.09 71.53 $\uparrow$ 74.37 74.69 $\uparrow$
maggic (n=20) 70.46 56.55 63.67 64.12 normal-$\uparrow$ 58.74 59.98 $\uparrow$ 53.27 55.37 $\uparrow$ 54.40 55.28 $\uparrow$ 50.39 55.27 $\uparrow$ 56.52 59.00 $\uparrow$ 55.83 56.45 $\uparrow$ 54.59 57.82 $\uparrow$
seer (n=20) 90.91 85.96 88.63 88.74 normal-$\uparrow$ 87.90 88.11 $\uparrow$ 78.22 83.69 $\uparrow$ 83.17 84.74 $\uparrow$ 51.07 77.84 $\uparrow$ 84.86 84.75 85.72 86.11 $\uparrow$ 80.33 82.63 $\uparrow$
compas (n=20) 73.02 63.74 71.88 71.10 66.62 67.89 $\uparrow$ 55.62 64.29 $\uparrow$ 62.15 64.70 $\uparrow$ 56.97 66.59 $\uparrow$ 64.71 68.59 $\uparrow$ 62.68 63.16 $\uparrow$ 61.19 64.36 $\uparrow$
adult (n=20) 88.37 79.24 45.28 71.07 $\uparrow$ 50.89 75.77 $\uparrow$ 80.03 81.27 normal-$\uparrow$ 73.19 74.10 $\uparrow$ 68.84 79.66 $\uparrow$ 75.65 80.74 $\uparrow$ 74.61 75.04 $\uparrow$ 76.82 78.84 $\uparrow$
drug (n=20) 81.41 73.56 78.54 78.10 71.69 71.68 73.07 74.88 $\uparrow$ 69.14 70.97 $\uparrow$ 62.38 69.99 $\uparrow$ 66.57 70.37 $\uparrow$ 71.99 72.28 $\uparrow$ 69.39 71.08 $\uparrow$
covid (n=40) 80.50 71.20 76.17 76.91 normal-$\uparrow$ 75.32 76.13 $\uparrow$ 64.56 68.82 $\uparrow$ 71.08 72.77 $\uparrow$ 61.40 71.21 $\uparrow$ 70.74 73.84 $\uparrow$ 68.85 69.65 $\uparrow$ 64.13 69.58 $\uparrow$
cutract (n=40) 79.14 71.90 74.05 74.82 normal-$\uparrow$ 72.69 72.89 $\uparrow$ 68.43 70.97 $\uparrow$ 70.67 72.39 $\uparrow$ 61.50 74.10 $\uparrow$ 67.98 71.55 $\uparrow$ 70.26 70.12 67.87 70.88 $\uparrow$
maggic (n=40) 70.36 59.65 61.89 63.39 normal-$\uparrow$ 60.41 61.84 $\uparrow$ 55.81 56.64 $\uparrow$ 56.58 58.08 $\uparrow$ 48.68 58.82 $\uparrow$ 57.50 60.90 $\uparrow$ 57.98 58.00 $\uparrow$ 56.61 59.42 $\uparrow$
seer (n=40) 90.93 87.86 88.50 88.89 normal-$\uparrow$ 85.66 85.62 85.95 87.62 $\uparrow$ 84.92 85.45 $\uparrow$ 60.75 85.84 $\uparrow$ 88.46 88.85 $\uparrow$ 86.56 87.58 $\uparrow$ 86.30 88.36 $\uparrow$
compas (n=40) 72.35 65.50 70.96 70.93 65.70 66.72 $\uparrow$ 60.24 63.78 $\uparrow$ 59.91 62.49 $\uparrow$ 65.17 68.19 $\uparrow$ 62.63 67.00 $\uparrow$ 63.70 64.18 $\uparrow$ 58.63 64.17 $\uparrow$
adult (n=40) 88.65 82.23 44.99 74.25 $\uparrow$ 56.38 76.76 $\uparrow$ 78.24 82.22 $\uparrow$ 71.83 81.15 $\uparrow$ 73.65 81.54 $\uparrow$ 81.49 84.34 normal-$\uparrow$ 81.80 83.16 $\uparrow$ 80.54 83.02 $\uparrow$
drug (n=40) 81.26 71.74 78.98 78.88 71.54 73.21 $\uparrow$ 71.34 74.44 $\uparrow$ 72.12 73.68 $\uparrow$ 70.00 76.45 $\uparrow$ 66.91 73.78 $\uparrow$ 69.09 69.75 $\uparrow$ 68.21 70.94 $\uparrow$
covid (n=100) 80.59 74.19 76.53 77.34 normal-$\uparrow$ 74.74 76.27 $\uparrow$ 72.30 75.15 $\uparrow$ 75.45 76.20 $\uparrow$ 69.05 74.94 $\uparrow$ 71.01 76.38 $\uparrow$ 71.71 72.28 $\uparrow$ 73.90 75.17 $\uparrow$
cutract (n=100) 78.93 76.73 74.62 75.70 $\uparrow$ 73.75 74.40 $\uparrow$ 73.89 75.61 $\uparrow$ 75.70 74.49 60.03 74.35 $\uparrow$ 75.49 76.86 normal-$\uparrow$ 76.13 75.47 73.85 76.21 $\uparrow$
maggic (n=100) 70.48 60.14 61.31 63.64 normal-$\uparrow$ 57.17 60.28 $\uparrow$ 59.32 61.13 $\uparrow$ 58.12 58.97 $\uparrow$ 49.10 59.95 $\uparrow$ 60.43 63.06 $\uparrow$ 59.87 59.59 59.43 61.28 $\uparrow$
seer (n=100) 90.93 88.92 88.16 88.69 $\uparrow$ 88.09 88.25 $\uparrow$ 88.45 89.11 $\uparrow$ 87.53 88.25 $\uparrow$ 83.13 88.44 $\uparrow$ 88.46 89.27 normal-$\uparrow$ 87.95 87.65 86.81 88.67 $\uparrow$
compas (n=100) 72.30 67.75 71.21 71.34 normal-$\uparrow$ 65.61 67.67 $\uparrow$ 64.14 66.72 $\uparrow$ 61.55 65.03 $\uparrow$ 67.43 68.88 $\uparrow$ 65.84 68.39 $\uparrow$ 66.39 66.53 $\uparrow$ 65.61 67.77 $\uparrow$
adult (n=100) 88.42 85.59 46.71 78.68 $\uparrow$ 48.64 75.94 $\uparrow$ 82.30 84.48 $\uparrow$ 77.41 78.98 $\uparrow$ 83.99 84.99 $\uparrow$ 82.67 85.77 normal-$\uparrow$ 85.63 85.43 81.32 83.97 $\uparrow$
drug (n=100) 81.42 74.34 80.35 80.01 70.00 71.76 $\uparrow$ 71.66 75.80 $\uparrow$ 71.26 72.40 $\uparrow$ 74.57 78.65 $\uparrow$ 69.34 77.50 $\uparrow$ 71.04 71.75 $\uparrow$ 71.91 75.38 $\uparrow$
covid (n=200) 80.62 76.73 76.06 77.25 $\uparrow$ 75.18 77.09 $\uparrow$ 76.11 77.54 $\uparrow$ 77.51 78.09 normal-$\uparrow$ 71.92 76.62 $\uparrow$ 74.06 77.53 $\uparrow$ 74.94 74.94 72.75 76.00 $\uparrow$
cutract (n=200) 78.58 77.53 75.28 76.30 $\uparrow$ 74.36 75.38 $\uparrow$ 75.25 75.88 $\uparrow$ 77.66 77.66 75.37 76.69 $\uparrow$ 76.45 77.50 $\uparrow$ 76.77 76.59 72.94 74.86 $\uparrow$
maggic (n=200) 70.28 63.32 61.89 64.30 normal-$\uparrow$ 59.10 62.10 $\uparrow$ 61.57 63.98 $\uparrow$ 56.77 57.35 $\uparrow$ 51.05 63.17 $\uparrow$ 59.17 63.84 $\uparrow$ 62.86 62.86 60.94 63.47 $\uparrow$
seer (n=200) 90.94 90.01 88.65 89.20 $\uparrow$ 87.57 88.14 $\uparrow$ 88.20 88.80 $\uparrow$ 89.69 89.81 $\uparrow$ 89.10 89.68 $\uparrow$ 87.35 89.54 $\uparrow$ 89.13 89.14 $\uparrow$ 88.74 89.65 $\uparrow$
compas (n=200) 72.47 69.14 70.53 71.36 normal-$\uparrow$ 63.49 66.50 $\uparrow$ 66.26 68.73 $\uparrow$ 64.83 66.40 $\uparrow$ 67.85 69.93 $\uparrow$ 60.78 68.07 $\uparrow$ 68.18 68.18 66.26 68.99 $\uparrow$
adult (n=200) 88.50 86.94 38.89 80.35 $\uparrow$ 56.97 79.07 $\uparrow$ 85.25 86.29 $\uparrow$ 86.81 86.63 85.70 85.94 $\uparrow$ 84.18 86.65 $\uparrow$ 86.88 86.70 84.09 85.91 $\uparrow$
drug (n=200) 81.33 77.47 79.06 79.61 $\uparrow$ 71.56 72.54 $\uparrow$ 76.15 78.84 $\uparrow$ 69.00 73.14 $\uparrow$ 77.51 79.72 $\uparrow$ 75.47 80.07 normal-$\uparrow$ 75.82 75.69 71.76 76.95 $\uparrow$
C.2 Full results for performance evaluation
We report full results with standard deviation for the results from the main paper. The performance is AUC averaged over XGBoost, Random forest, Logistic regression, Decision tree.
Table 10: AUC averaged over 4 downstream models on $D_{\mathrm{test}}$ where curation improves performance for all methods across all sample sizes $n$, as indicated by $\uparrow$. CLLM  w/ GPT-4 (Curated) dataset provides the strongest performance for both private/proprietary datasets and public datasets
Real data CLLM (OURS) Baselines
GPT-4 GPT-3.5 CTGAN TabDDPM GReaT NFLOW SMOTE TVAE
Dataset $D_{\mathrm{oracle}}$ $D_{\mathrm{train}}$ Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur. Uncur. Cur.
covid (n=20) $74.41_{(0.11)}$ $68.50_{(1.57)}$ $73.78_{(0.31)}$ $\textbf{73.87}_{(0.50)}$ $69.85_{(0.75)}$ $71.41_{(0.92)}$ $59.00_{(2.25)}$ $63.67_{(2.51)}$ $66.84_{(1.66)}$ $66.85_{(1.56)}$ $57.38_{(1.47)}$ $66.46_{(0.80)}$ $62.87_{(0.98)}$ $68.56_{(1.07)}$ $66.95_{(1.66)}$ $66.82_{(1.89)}$ $61.69_{(2.72)}$ $66.11_{(2.79)}$
cutract (n=20) $72.23_{(0.65)}$ $70.12_{(1.16)}$ $71.15_{(0.46)}$ $\textbf{72.50}_{(0.76)}$ $69.97_{(1.13)}$ $71.54_{(1.38)}$ $64.01_{(2.26)}$ $67.98_{(1.63)}$ $66.05_{(1.14)}$ $66.59_{(1.49)}$ $52.38_{(1.78)}$ $67.02_{(0.97)}$ $64.44_{(1.05)}$ $70.42_{(1.25)}$ $68.41_{(1.34)}$ $69.24_{(1.25)}$ $68.94_{(1.38)}$ $70.22_{(1.12)}$
maggic (n=20) $67.41_{(0.06)}$ $57.13_{(0.85)}$ $60.70_{(0.38)}$ $\textbf{61.48}_{(0.49)}$ $57.54_{(0.83)}$ $58.69_{(0.76)}$ $52.75_{(1.34)}$ $54.51_{(1.40)}$ $54.59_{(1.02)}$ $55.39_{(1.07)}$ $50.29_{(0.52)}$ $55.64_{(0.43)}$ $54.72_{(1.38)}$ $57.38_{(1.06)}$ $55.84_{(1.21)}$ $56.15_{(1.20)}$ $54.08_{(1.05)}$ $56.19_{(0.76)}$
seer (n=20) $87.92_{(0.08)}$ $80.67_{(1.67)}$ $84.53_{(0.33)}$ $\textbf{84.82}_{(0.47)}$ $83.34_{(0.95)}$ $83.71_{(0.52)}$ $74.34_{(4.11)}$ $78.73_{(3.13)}$ $80.59_{(1.32)}$ $80.60_{(1.32)}$ $47.57_{(2.51)}$ $74.43_{(3.36)}$ $76.06_{(1.68)}$ $79.98_{(1.65)}$ $79.23_{(2.17)}$ $80.02_{(2.08)}$ $74.53_{(1.37)}$ $78.73_{(1.24)}$
compas (n=20) $67.51_{(0.03)}$ $63.11_{(0.96)}$ $\textbf{68.01}_{(0.44)}$ $67.91_{(0.55)}$ $62.07_{(1.51)}$ $64.43_{(1.38)}$ $55.67_{(1.60)}$ $62.56_{(1.61)}$ $57.67_{(1.90)}$ $60.87_{(1.33)}$ $53.33_{(1.85)}$ $63.59_{(1.30)}$ $59.49_{(1.72)}$ $64.62_{(1.14)}$ $61.06_{(1.92)}$ $61.59_{(2.15)}$ $58.30_{(1.77)}$ $62.58_{(1.87)}$
adult (n=20) $84.17_{(0.10)}$ $77.45_{(1.25)}$ $50.39_{(3.99)}$ $71.48_{(2.34)}$ $49.23_{(2.57)}$ $72.37_{(2.26)}$ $72.23_{(1.26)}$ $76.86_{(1.25)}$ $74.35_{(1.48)}$ $75.04_{(1.82)}$ $67.00_{(4.72)}$ $77.25_{(1.49)}$ $67.46_{(3.76)}$ $76.48_{(1.77)}$ $73.75_{(1.64)}$ $73.67_{(1.46)}$ $73.20_{(1.51)}$ $76.90_{(1.64)}$
drug (n=20) $77.81_{(0.55)}$ $70.84_{(2.25)}$ $75.08_{(1.17)}$ $\textbf{75.29}_{(1.11)}$ $71.68_{(2.25)}$ $72.14_{(2.64)}$ $68.31_{(2.81)}$ $72.65_{(2.00)}$ $68.12_{(2.38)}$ $69.68_{(2.41)}$ $58.78_{(4.26)}$ $68.89_{(3.57)}$ $62.13_{(4.94)}$ $67.75_{(3.72)}$ $70.16_{(1.87)}$ $70.16_{(1.75)}$ $66.60_{(2.95)}$ $69.18_{(2.59)}$
covid (n=40) $75.02_{(0.22)}$ $70.77_{(0.96)}$ $73.40_{(0.61)}$ $\textbf{73.95}_{(0.67)}$ $70.42_{(0.92)}$ $71.93_{(0.60)}$ $63.63_{(1.15)}$ $68.46_{(0.93)}$ $70.50_{(1.49)}$ $70.44_{(1.37)}$ $56.50_{(0.83)}$ $68.68_{(1.39)}$ $66.41_{(2.51)}$ $70.48_{(1.48)}$ $68.66_{(1.48)}$ $68.44_{(1.28)}$ $61.03_{(0.96)}$ $67.35_{(0.60)}$
cutract (n=40) $72.57_{(0.48)}$ $69.18_{(0.65)}$ $69.87_{(0.62)}$ $\textbf{71.72}_{(0.46)}$ $68.47_{(0.59)}$ $69.56_{(0.50)}$ $63.01_{(2.33)}$ $67.87_{(1.36)}$ $65.63_{(2.71)}$ $67.27_{(2.38)}$ $54.39_{(1.59)}$ $68.44_{(0.41)}$ $61.40_{(1.95)}$ $67.98_{(1.06)}$ $67.86_{(0.72)}$ $67.95_{(0.59)}$ $59.79_{(1.85)}$ $66.62_{(1.06)}$
maggic (n=40) $67.50_{(0.04)}$ $58.26_{(0.55)}$ $59.29_{(0.50)}$ $\textbf{60.77}_{(0.36)}$ $57.50_{(0.69)}$ $59.15_{(0.48)}$ $55.00_{(1.27)}$ $56.78_{(1.17)}$ $55.24_{(0.63)}$ $56.94_{(0.50)}$ $48.81_{(0.73)}$ $56.64_{(0.63)}$ $54.68_{(0.53)}$ $58.58_{(0.54)}$ $57.40_{(0.57)}$ $57.44_{(0.67)}$ $55.04_{(1.14)}$ $57.33_{(1.06)}$
seer (n=40) $87.90_{(0.09)}$ $82.93_{(0.55)}$ $84.29_{(0.39)}$ $\textbf{84.93}_{(0.46)}$ $83.46_{(1.20)}$ $84.44_{(0.69)}$ $80.05_{(0.93)}$ $83.67_{(0.51)}$ $82.59_{(1.48)}$ $81.37_{(1.11)}$ $54.93_{(2.06)}$ $81.11_{(1.24)}$ $79.88_{(0.53)}$ $84.36_{(0.62)}$ $80.79_{(0.67)}$ $82.21_{(0.67)}$ $78.69_{(2.32)}$ $83.62_{(1.06)}$
compas (n=40) $67.35_{(0.16)}$ $62.34_{(0.79)}$ $67.57_{(0.40)}$ $\textbf{67.85}_{(0.37)}$ $61.34_{(1.67)}$ $62.84_{(1.21)}$ $56.29_{(1.96)}$ $61.02_{(1.69)}$ $58.85_{(1.28)}$ $60.11_{(1.22)}$ $58.88_{(1.04)}$ $64.37_{(0.91)}$ $58.61_{(1.35)}$ $63.54_{(1.11)}$ $60.83_{(1.19)}$ $60.95_{(1.22)}$ $55.94_{(1.50)}$ $61.04_{(1.56)}$
adult (n=40) $84.43_{(0.06)}$ $79.44_{(1.03)}$ $48.31_{(3.58)}$ $73.82_{(2.19)}$ $49.21_{(1.78)}$ $74.27_{(1.49)}$ $71.82_{(1.22)}$ $79.11_{(0.71)}$ $71.51_{(1.85)}$ $77.99_{(0.65)}$ $66.77_{(3.32)}$ $78.81_{(1.36)}$ $71.13_{(2.27)}$ $79.71_{(1.02)}$ $77.90_{(1.07)}$ $78.84_{(1.19)}$ $72.58_{(1.34)}$ $\textbf{80.02}_{(0.76)}$
drug (n=40) $77.71_{(0.32)}$ $71.86_{(1.07)}$ $74.30_{(0.59)}$ $\textbf{75.79}_{(0.39)}$ $71.33_{(0.88)}$ $72.76_{(0.97)}$ $69.46_{(2.23)}$ $72.74_{(1.70)}$ $71.08_{(1.72)}$ $73.07_{(1.03)}$ $64.89_{(1.39)}$ $73.64_{(0.87)}$ $62.51_{(3.03)}$ $70.97_{(1.91)}$ $69.23_{(1.68)}$ $69.78_{(1.46)}$ $65.22_{(1.37)}$ $70.30_{(1.04)}$
covid (n=100) $74.52_{(0.16)}$ $71.57_{(0.48)}$ $73.77_{(0.27)}$ $\textbf{74.71}_{(0.34)}$ $70.71_{(0.46)}$ $72.76_{(0.44)}$ $69.05_{(0.96)}$ $72.13_{(0.66)}$ $71.60_{(0.59)}$ $73.22_{(0.46)}$ $63.52_{(1.29)}$ $72.04_{(0.57)}$ $64.25_{(1.51)}$ $72.64_{(0.64)}$ $70.08_{(0.67)}$ $70.78_{(0.59)}$ $69.05_{(0.48)}$ $71.96_{(0.49)}$
cutract (n=100) $72.36_{(0.47)}$ $70.96_{(0.68)}$ $70.20_{(0.45)}$ $\textbf{72.51}_{(0.55)}$ $69.97_{(0.97)}$ $71.94_{(0.89)}$ $67.94_{(1.01)}$ $72.42_{(0.66)}$ $70.53_{(1.51)}$ $71.98_{(1.39)}$ $55.72_{(2.04)}$ $69.14_{(0.93)}$ $67.59_{(0.71)}$ $72.42_{(0.54)}$ $68.79_{(0.83)}$ $69.68_{(0.76)}$ $66.89_{(1.03)}$ $71.52_{(0.70)}$
maggic (n=100) $67.46_{(0.07)}$ $59.65_{(0.50)}$ $58.98_{(0.23)}$ $\textbf{61.32}_{(0.42)}$ $55.71_{(0.83)}$ $58.90_{(0.72)}$ $57.20_{(0.91)}$ $59.34_{(0.64)}$ $57.26_{(0.50)}$ $58.28_{(0.46)}$ $49.54_{(0.71)}$ $57.91_{(0.74)}$ $56.36_{(0.54)}$ $60.11_{(0.54)}$ $58.89_{(0.51)}$ $58.99_{(0.42)}$ $56.17_{(0.68)}$ $58.86_{(0.69)}$
seer (n=100) $87.79_{(0.07)}$ $83.95_{(0.32)}$ $84.45_{(0.38)}$ $\textbf{85.37}_{(0.47)}$ $83.92_{(0.81)}$ $85.08_{(0.32)}$ $81.60_{(0.73)}$ $85.14_{(0.36)}$ $83.04_{(0.78)}$ $84.83_{(0.49)}$ $70.32_{(2.52)}$ $83.83_{(0.39)}$ $81.16_{(1.03)}$ $85.03_{(0.39)}$ $81.82_{(0.45)}$ $82.49_{(0.46)}$ $78.88_{(0.71)}$ $84.50_{(0.44)}$
compas (n=100) $67.18_{(0.30)}$ $62.56_{(0.72)}$ $68.02_{(0.29)}$ $\textbf{68.19}_{(0.37)}$ $60.10_{(1.60)}$ $62.47_{(1.11)}$ $60.01_{(1.16)}$ $63.73_{(0.99)}$ $58.32_{(0.91)}$ $61.34_{(0.94)}$ $59.97_{(0.82)}$ $64.19_{(0.83)}$ $60.02_{(0.63)}$ $64.04_{(0.60)}$ $61.44_{(0.84)}$ $61.73_{(0.81)}$ $59.97_{(1.00)}$ $62.82_{(1.05)}$
adult (n=100) $84.34_{(0.07)}$ $81.24_{(0.48)}$ $46.09_{(1.86)}$ $74.57_{(1.74)}$ $47.56_{(3.43)}$ $73.97_{(1.57)}$ $74.29_{(1.23)}$ $80.45_{(0.81)}$ $75.93_{(1.45)}$ $78.22_{(1.22)}$ $77.09_{(0.74)}$ $\textbf{81.66}_{(0.53)}$ $70.70_{(}{1.02)}$ $81.04_{(0.37)}$ $80.56_{(0.53)}$ $81.10_{(0.44)}$ $74.04_{(1.44)}$ $80.23_{(0.98)}$
drug (n=100) $78.00_{(0.19)}$ $73.58_{(0.72)}$ $76.24_{(0.64)}$ $\textbf{76.74}_{(0.51)}$ $69.46_{(2.43)}$ $71.05_{(2.30)}$ $68.19_{(2.06)}$ $73.28_{(1.41)}$ $72.43_{(0.96)}$ $73.79_{(0.65)}$ $67.26_{(1.22)}$ $75.28_{(0.46)}$ $62.67_{(2.23)}$ $73.12_{(0.97)}$ $70.90_{(1.08)}$ $71.53_{(1.13)}$ $68.22_{(1.60)}$ $73.59_{(0.97)}$
covid (n=200) $74.69_{(0.19)}$ $72.33_{(0.52)}$ $73.40_{(0.25)}$ $\textbf{74.62}_{(0.13)}$ $70.70_{(0.73)}$ $73.12_{(0.48)}$ $71.07_{(0.35)}$ $73.89_{(0.40)}$ $72.47_{(0.50)}$ $74.44_{(0.44)}$ $65.55_{(0.67)}$ $73.07_{(0.38)}$ $65.04_{(0.69)}$ $72.90_{(0.37)}$ $71.68_{(0.49)}$ $71.87_{(0.49)}$ $67.89_{(0.56)}$ $72.38_{(0.42)}$
cutract (n=200) $72.52_{(0.49)}$ $71.75_{(0.71)}$ $71.39_{(0.76)}$ $73.01_{(0.85)}$ $70.28_{(0.67)}$ $72.39_{(0.84)}$ $69.28_{(0.55)}$ $72.41_{(0.65)}$ $71.83_{(0.65)}$ $\textbf{74.03}_{(0.69)}$ $66.66_{(0.94)}$ $72.49_{(0.71)}$ $68.77_{(0.71)}$ $73.16_{(0.68)}$ $70.23_{(0.83)}$ $70.80_{(0.79)}$ $66.61_{(0.76)}$ $71.87_{(0.71)}$
maggic (n=200) $67.37_{(0.06)}$ $61.39_{(0.44)}$ $58.92_{(0.47)}$ $\textbf{61.41}_{(0.39)}$ $57.33_{(0.51)}$ $60.16_{(0.41)}$ $58.48_{(0.59)}$ $61.33_{(0.57)}$ $56.26_{(0.82)}$ $57.20_{(0.91)}$ $50.74_{(0.74)}$ $59.60_{(0.55)}$ $55.95_{(0.72)}$ $60.75_{(0.39)}$ $60.73_{(0.44)}$ $60.78_{(0.49)}$ $57.18_{(0.56)}$ $60.23_{(0.44)}$
seer (n=200) $87.84_{(0.08)}$ $84.63_{(0.35)}$ $84.39_{(0.19)}$ $85.56_{(0.30)}$ $83.48_{(0.41)}$ $84.80_{(0.34)}$ $82.04_{(0.70)}$ $85.34_{(0.36)}$ $84.39_{(0.44)}$ $\textbf{86.57}_{(0.27)}$ $82.15_{(0.43)}$ $86.03_{(0.23)}$ $77.73_{(1.65)}$ $85.19_{(0.22)}$ $83.38_{(0.32)}$ $84.15_{(0.42)}$ $79.71_{(0.58)}$ $85.26_{(0.25)}$
compas (n=200) $67.14_{(0.20)}$ $63.27_{(0.60)}$ $67.02_{(0.53)}$ $\textbf{68.15}_{(0.40)}$ $60.48_{(1.86)}$ $63.39_{(1.28)}$ $60.58_{(0.83)}$ $64.32_{(0.78)}$ $60.60_{(0.90)}$ $63.52_{(1.04)}$ $61.11_{(0.63)}$ $65.08_{(0.53)}$ $56.58_{(1.03)}$ $63.60_{(0.83)}$ $61.99_{(0.64)}$ $62.80_{(0.61)}$ $60.15_{(1.14)}$ $63.99_{(0.71)}$
adult (n=200) $84.25_{(0.04)}$ $82.12_{(0.41)}$ $40.96_{(2.47)}$ $75.84_{(1.43)}$ $49.89_{(3.22)}$ $76.11_{(1.43)}$ $78.18_{(0.26)}$ $82.32_{(0.29)}$ $81.66_{(0.19)}$ $83.17_{(0.20)}$ $80.06_{(0.53)}$ $\textbf{83.32}_{(0.35)}$ $74.31_{(0.81)}$ $82.64_{(0.33)}$ $82.26_{(0.38)}$ $82.39_{(0.33)}$ $75.21_{(0.69)}$ $82.02_{(0.25)}$
drug (n=200) $77.36_{(0.51)}$ $76.10_{(0.45)}$ $75.58_{(0.55)}$ $76.06_{(0.42)}$ $70.66_{(1.56)}$ $72.81_{(1.18)}$ $71.31_{(1.17)}$ $75.98_{(0.76)}$ $69.61_{(2.76)}$ $71.79_{(1.99)}$ $72.35_{(0.61)}$ $\textbf{77.41}_{(0.56)}$ $65.25_{(2.47)}$ $75.26_{(0.77)}$ $74.38_{(0.50)}$ $74.78_{(0.54)}$ $68.39_{(1.26)}$ $74.33_{(0.61)}$
C.3 Decoupling prior knowledge and data model
Two components can be attributed to the good performances of CLLM: the background knowledge of the LLM, and its capacity to build a strong data model. In this subsection, we provide insights to understand the effect of the LLM’s background knowledge (e.g. prior).
Table 11: Downstream accuracy when varying the number of in-context samples in the prompt to generate the augmented datasets.
In-context samples Downstream accuracy
$n=1$ (Prior) $70.20\pm 1.60$
$n=20$ $73.87\pm 0.50$
$n=40$ $73.95\pm 0.67$
$n=100$ $74.71\pm 0.34$
$D_{\mathrm{oracle}}$ $74.6\pm 0.15$
We considered the Covid dataset (private medical dataset, to avoid memorization issues) and generated data with GPT-4 (same as Section 2.1). We ablate the prompt used in our work (detailed in Appendix B.5), and solely provide one in-context example in the prompt, in order to give the LLM the minimal amount of information about the desired structure of the dataset. This forces the LLM to rely on its own prior (background knowledge), and removes the effect of in-context examples which could be used to build a data model. We report the results for the prior and CLLM in Table 11.
From these results, we conclude that the LLM prior permits to obtain good downstream performance, but is outperformed by $\mathcal{D_{\mathrm{oracle}}}$ by a margin of $4.4\%$. Hence, we cannot solely rely on the prior. Furthermore, downstream performance increases as the number of in-context samples increases. This shows it is indeed important to include the in-context examples if we wish to obtain downstream performance close to $\mathcal{D}_{\mathrm{oracle}}$, as the LLM can build a good data model. This implies that while the LLM does use background knowledge of similar datasets, it still requires in-context samples to refine its prior by creating a good data model.
We then quantify and visualize the strength of the prior, by studying how much the LLMs output distribution adapts to the in-context samples provided. We evaluate data generated by the prior of the LLM ($n=1$), and for $n=20,40,100$ on the Covid dataset.
In particular, we observe in Figure 8 that there is a region in the oracle data which is not captured by the LLM’s prior output (the left part of the leftmost blob, circled in blue in Figure 8). However, as the number of in-context real examples increases in the prompt of the LLM, we observe that this steers the LLM to generate data which covers this region. This region is associated to the subgroup of people older than $87$ years old, and having many severe comorbidities (e.g. Diabetes, Cardiovascular diseases) and many respiratory symptoms. This subgroup, in the Oracle dataset, represents less than $3.5\%$ of the data, and is completely ignored by the GPT-4 prior. In particular, the prior defaults to more typical patients in the range $70$-$80$ years old. On the contrary, as $n$ increases, the LLM is guided by the in-context samples and generates samples from this subgroup, which are ”rarer” or different from the general population.
This demonstrates that the LLM captures the distinct features of this particular region, and hence is not overwhelmed by the prior, but instead the data in the form of in-context samples adapts it, hence aligning the augmented dataset with the ground-truth distribution.
Figure 8: The data generated by the LLM captures the distinct features in ”atypical” regions of the Oracle manifold, as in-context samples are added to the prompt. This shows that it is flexible enough to adapt its prior knowledge to the nuances of the data. The group encircled in blue represents patients who are $>88$ years old, representing around $3.5\%$ of the Oracle. This illustrates the added in-context samples can successfully guide the LLM to generate these rare samples.
C.4 Ablation for contextual information on Compas
We conduct a similar experiment as in Table 2.1, and use the dataset Compas. We report the results in Table C.4.
Table 12: Including contextual information in the prompt improves precision (P), recall (R), and utility (U) in low-sample settings (results shown for Compas).
$n_{\mathrm{samples}}$ in $D_{\mathrm{train}}$ GPT-4 w/ context GPT-4 no context TVAE
P R U P R U P R U
20 $\bf 0.69_{(0.02)}$ $\bf 0.88_{(0.02)}$ $\bf 0.69_{(0.02)}$ $0.27_{(0.03)}$ $\bf 0.89_{(0.03)}$ $0.60_{(0.03)}$ $0.43_{(0.02)}$ $0.43_{(0.05)}$ $0.55_{(0.04)}$
40 $\bf 0.70_{(0.0)}$ $\bf 0.92_{(0.01)}$ $\bf 0.65_{(0.03)}$ $0.31_{(0.06)}$ $0.84_{(0.03)}$ $0.57_{(0.01)}$ $0.54_{(0.02)}$ $0.80_{(0.02)}$ $0.50_{(0.04)}$
100 $\bf 0.69_{(0.02)}$ $\bf 0.89_{(0.02)}$ $\bf 0.69_{(0.01)}$ $0.34_{(0.1)}$ $0.85_{(0.05)}$ $0.62_{(0.01)}$ $0.60_{(0.03)}$ $0.86_{(0.02)}$ $0.59_{(0.03)}$
200 $\bf 0.70_{(0.01)}$ $\bf 0.89_{(0.02)}$ $\bf 0.69_{(0.01)}$ $0.31_{(0.05)}$ $0.87_{(0.03)}$ $0.58_{(0.05)}$ $0.65_{(0.02)}$ $0.88_{(0.01)}$ $0.63_{(0.01)}$
These results highlight the importance of incorporating contextual information in the prompt, as it enables to exploit the prior knowledge of the LLM.
C.5 Comparison to random noise baseline
We compare CLLM to a random noise baseline, where we augment the dataset with random additive Gaussian noise. In order to capture the correlations between the different features, we fit a Kernel Density Estimator with a Gaussian kernel and bandwidth given by Scott’s rule. We then sample $1000$ points to create an augmented dataset $D_{\mathrm{syn}}$. We report the performance gap between CLLM and this baseline (with and without curation) for the Covid and Compas datasets in Figure 9. We observe that the random noise baseline does not match the performance of CLLM (i.e. has a performance gap), although the baseline naturally improves as the dataset $D_{\mathrm{train}}$ grows in size.
Figure 9: The random noise baseline does not match the performance of CLLM